### v4

# Steering Flexible Linear Objects in Planar Environments by Two Robot Hands Using Euler’s Elastica Solutions

Abstract
The manipulation of flexible objects such as cables, wires and fresh food items by robot hands forms a special challenge in robot grasp mechanics. This paper considers the steering of flexible linear objects in planar environments by two robot hands. The flexible linear object, modeled as an elastic non-stretchable rod, is manipulated by varying the gripping endpoint positions while keeping equal endpoint tangents. The flexible linear object shape has a closed form solution in terms of the grasp endpoint positions and tangents, called Euler’s elastica. This paper obtains the elastica solutions under the optimal control framework, then uses the elastica solutions to obtain closed form criteria for non self-intersection, stability and obstacle avoidance of the flexible linear object. The new tools are incorporated into a planning scheme for steering flexible linear objects in planar environments populated by sparsely spaced obstacles. The scheme is fully implemented and demonstrated with detailed examples.
I Introduction
Robotic manipulation of flexible linear objects such as cables, wires and fresh food items forms a special challenge in robot grasp mechanics. In these problems, one or two robot hands apply endpoint forces and torques that together with external influences such as gravity and contacts with the environment affect the object shape during manipulation. Robotic applications include cable routing and untangling [cable_routing11, bretl_multiple, cable_rss22], surgical suturing [jackson&cavusogl, javdani], knot tying [balkcom15, hirai_knot], compliant mechanisms [till17], fresh food handling [food_survey22], architectural elements fabrication [brander] and agricultural robotics [adaptive_LDO22].
This paper focuses on robotic steering of flexible linear objects in planar environments. The object, modeled as a non-stretchable elastic rod, is to be steered from start to target positions by two robot hands while avoiding self-collision and contact with the environment, except for endpoint contacts at the start or target positions (Fig. 1). Examples of flexible linear objects that can be steered in two-dimensions are strip like objects such as ribbon cables, plastic ties and fresh food items, since such objects preserve their strip flatness during manipulation by two robot hands. This paper provides analysis of flexible linear objects mechanics in two-dimensions, then a scheme for steering such objects based on closed form solutions for their shape and stability that depend only on the gripping hands relative position during manipulation.
Figure 1: Top view of a flexible cable steered by two robot grippers in a planar environment populated by obstacles. The cable must reach target position while avoiding self-collision and contact with obstacles (except for endpoint contacts at the start or target).
Related work: We use the term flexible cable for flexible linear objects manipulated in two-dimensions. The modeling of these objects relies on Euler-Bernoulli bending moment law [goss_thesis]. When a flexible cable is held at an equilibrium state by endpoint forces and moments in the plane, the cable curvature at every point is proportional to the cable bending moment at this point. The coefficient of proportionality is the cable stiffness at each point.
The mechanics literature has used the Euler-Bernoulli law to obtain solutions for flexible cable equilibrium shapes in two-dimensions [love]. The solutions, called Euler’s elastica, form a one-parameter family of periodic shapes aligned along a linear axis [levyakov10]. The axis is parallel to the opposing forces applied at the cable endpoints and passes through the zero curvature cable points. While flexible cable equilibrium shapes have been fully characterized in the mechanics literature, verification of their stability typically relies on numerical techniques [batista15, levyakov09]. Sachkov [sachkov07, sachkov_conjugate, sachkov_vertices] describes analytic bounds for flexible cable stability that are used in this paper’s steering scheme.
In the robotics literature, sampling based approaches are used to plan flexible cable steering paths. Moll [kavraki06] samples cable endpoint positions, then computes their stable shapes by numerical optimization of the flexible cable total elastic energy. Bretl [bretl_ijrr14] uses Sachkov’s approach to describe the cable total elastic energy minimization as an optimal control problem. Bretl shows that the adjoint differential equation which describes the equilibrium shapes of a flexible cable held by robot hands is fully determined by a small number of co-state variables [bretl_ijrr14]. In 2-D environments these are endpoint forces and moments (three variables) while in 3-D environments these are endpoint forces and toques (six variables).
Using this insight, Bretl developed sampling based planners for steering flexible cables in 2-D and 3-D environments [bretl_ijrr14]. However, each sampled co-state requires solution of the adjoint differential equation to obtain the cable shape, then computation of the cable co-state to endpoint Jacobian to verify that each sampled cable shape is stable. Sintov [sintov20] extended this work into a two stage approach. First a roadmap of stable cable shapes is computed for sampled co-states. Then a cable steering path is computed in the physical environment using numerical self-intersection and obstacle avoidance tests.
While this paper focuses on Euler’s elastica as a means to model flexible object shapes, robotic manipulation of heterogeneous flexible objects in agricultural robotics [zhu_survey22] and fresh food handling [food_survey22] requires complementary approaches. Papers such as [adaptive_LDO22], [zhu_iros18] and [lagneau20] describe adaptive and learning based approaches to manipulation of such complex flexible objects. One hopes that this paper’s modeling approach will become part of adaptive and learning based manipulation techniques for complex flexible objects.
Paper contributions: This paper considers steering flexible cables in two-dimensions using two robot hands. By maintaining equal endpoint tangents, the cable steering problem is reduced to five configuration variables: the cable base-frame configuration and the cable endpoints relative position. The paper starts with a derivation of Euler’s elastica solutions for flexible cable equilibrium shapes under the optimal control approach. The derivation complements Bretle [bretl_ijrr14] that focused on the adjoint equation and the use of its co-states to steer flexible cables in 3-D environments. This paper focuses on flexible cable steering in two-dimensions in order to take advantage of the closed form elastica solutions.
When considering flexible cables in 2-D environments, Euler’s elastica form periodic shapes aligned with a linear axis. This paper describes analytic equations that determine the cable shape in terms of the cable’s relative endpoint positions imposed by the robot hands. The paper next describes the range of the elastica modulus parameter that ensures non self-intersection, followed by a simple geometric rule that ensures stability of the flexible cable during steering.
To ensure obstacle avoidance during steering, the paper describes an approximation of the flexible cable equilibrium shapes by quadratic arcs determined by control points located on the cable. The piecewise quadratic approximation allows efficient collision checks against obstacles in the environment. The new tools are incorporated into a fully implemented planning scheme for steering flexible linear objects in planar environments populated by sparsely spaced obstacles.
The paper is structured as follows. Section II derives the flexible cable equilibrium shapes and formulates the flexible cable steering problem. Section III characterizes the non self-intersecting flexible cable shapes. Section IV describes a geometric rule that ensures flexible cable stability during steering. Section V describes an approximation of the flexible cable equilibrium shapes by quadratic arcs. Section VI incorporates these tools into a motion planning scheme that steers flexible cable among sparsely spaced obstacles. Section VII describes execution examples of the steering scheme. The conclusion discusses future research such as gravitational effects and interaction with the environment. An appendix verifies formulas for flexible cable curvature and tangent used in this paper.
II Euler’s Elastica as an Optimal Control Problem
This section obtains the flexible cable equilibrium shapes as an optimal control problem solution. The solution parameters are then related to the cable endpoint positions and the cable steering problem is formulated.
Consider a non-stretchable flexible cable of length $L$ in $\mbox{$I\!\!R$}^{2}$. The cable is parametrized by $(x\mbox{\small$(s)$},y\mbox{\small$(s)$})$ for $s\!\in\![0,L]$. The cable state variables are its $(x\mbox{\small$(s)$},y\mbox{\small$(s)$})$ coordinates and tangent direction $\phi(s)$, thus defining the state vector $S\!=\!(x,y,\phi)$. Under arclength parametrization with unit norm tangent, the cable curvature is given by $\mbox{$\kappa$}(s)\!=\!\tfrac{d}{ds}\phi(s)$. The cable curvature forms a continuous and piecewise smooth control input, $u(s)\!=\!\mbox{$\kappa$}(s)$, for the cable system equations given by
$\begin{array}[]{cl}\dot{x}(s)&=\cos\phi(s)\\ \dot{y}(s)&=\sin\phi(s)\\ \dot{\phi}(s)&=u(s)\end{array}\hskip 10.00002pt\mbox{\hskip 15.00002pt $s\in[0,L]$.}\vskip-5.05942pt$ (1)
Eq. (1) states that the cable shape, parametrized by arclength, is determined by its curvature as control input. When the flexible cable is modeled as an elastic rod and there are no external influences such as gravity and contacts with the environment, the cable’s total elastic energy is given by
$\mbox{$\cal E$}=\int_{0}^{L}\!\mbox{$\frac{1}{2}$}\mbox{\small$EI$}\!\cdot\!\mbox{$\kappa$}^{2}(s)ds\vskip-4.33601pt$ (2)
where E>0E\!>\!0 is the cable Young’s modulus of elasticity and I>0I\!>\!0 is the cable cross-sectional 2’nd moment of inertia [timoshenko&goodier:1969]. The cable stiffness, $EI$, is assumed to be a known parameter.
[…]
$\begin{array}[]{cl}&H(\mbox{\small$S$}(s),\mbox{$\lambda$}(s),u(s))\\ &=\mbox{$\lambda$}_{x}(s)\cos\phi(s)\!+\!\mbox{$\lambda$}_{y}(s)\sin\phi(s)\!+\!\mbox{$\lambda$}_{\phi}(s)u(s)\!+\!\mbox{$\frac{1}{2}$}\mbox{$\scriptsize EI$}\!\cdot\!u^{2}(s)\end{array}\vskip-3.61371pt$
where $\mbox{$\lambda$}(s)\!=\!(\mbox{$\lambda$}_{x}(s),\mbox{$\lambda$}_{y}(s),\mbox{$\lambda$}_{\phi}(s))$ are the co-state variables. The co-states $(\mbox{$\lambda$}_{x}(s),\mbox{$\lambda$}_{y}(s))$ and $\mbox{$\lambda$}_{\phi}(s)$ represent internal force and bending moment that develop at the cable points. The energy extremal cable shapes represent equilibrium states of the cable held by the robot hands. Along an extremal cable shape, the vector $\mbox{$\lambda$}(s)$ is determined by the adjoint equation [pontryagin]
$\dot{\lambda}(s)=-\frac{\partial}{\partial S}H\big(\mbox{\small$S$}(s),\mbox{$\lambda$}(s),u(s)\big)\hskip 10.00002pt\mbox{\hskip 10.00002pt\mbox{$S=(x,y,\phi)$}}\vskip-2.8903pt$
thus leading to the system of adjoint differntial equations
$\begin{split}\dot{\mbox{$\lambda$}}_{x}(s)&=0\\ \dot{\mbox{$\lambda$}}_{y}(s)&=0\\ \dot{\mbox{$\lambda$}}_{\phi}(s)&=\mbox{$\lambda$}_{x}(s)\sin\phi(s)\!-\!\mbox{$\lambda$}_{y}(s)\cos\phi(s)\end{split}$ (3)
while the control $u(s)\!=\!\mbox{$\kappa$}(s)$ satisfies the additional condition
$\frac{\partial}{\partial u}H\big(\mbox{\small$S$}(s),\mbox{$\lambda$}(s),u(s)\big)=0,\vskip-4.33601pt$ (4)
thus leading to the algebraic equation
$\mbox{$\lambda$}_{\phi}(s)+\mbox{\small$EI$}\!\cdot\!u(s)=0\vskip-6.50403pt$ (5)
which is the Euler-Bernoulli bending moment law.
Figure 2: Top view of a flexible cable of length $L$ embedded in its periodic elastica solution. The elastica axis with angle $\phi_{0}$ passes through the elastica zero curvature points and is parallel to the opposing forces of magnitude $\mbox{$\lambda$}_{r}$ applied at the cable endpoints.
From Eq. (3) one obtains that $\mbox{$\lambda$}_{x}(s)$ and $\mbox{$\lambda$}_{y}(s)$ are constants along energy extremal cable shapes. These constants define the co-state parameters $\mbox{\small$\lambda$}_{r}$ and $\phi_{0}$
[…]
which represent the magnitude and direction of the opposing forces applied at the cable endpoints (blue arrows in Fig. 2). The system described by Eq. (1) is autonomous (no terms explicitly depend on $s$). Hence the Hamiltonian is constant along energy extremal cable shapes, $H(s)\!=\!H^{*}$ for $s\!\in\![0,L]$. Substituting $\mbox{\small$\lambda$}_{\phi}(s)\!=\!-\mbox{\small$EI$}\!\cdot\!u(s)$ according to Eq. (5) in $H(s)$ and then substituting $u(s)\!=\!\mbox{$\kappa$}(s)$ gives
$\mbox{\small$\lambda$}_{r}\cdot\big(\cos\phi\mbox{\small$(s)$}\cos\phi_{0}+\sin\phi\mbox{\small$(s)$}\sin\phi_{0}\big)-\mbox{$\frac{1}{2}$}\mbox{\small$EI$}\cdot\mbox{$\kappa$}^{2}(s)=H^{*}.\vskip-5.05942pt$ (6)
Taking the derivative of both sides with respect to $s$ gives
$\mbox{$\lambda$}_{r}\!\cdot\!\big(-\!\sin\phi\mbox{\small$(s)$}\cos\phi_{0}+\cos\phi\mbox{\small$(s)$}\sin\phi_{0}\big)-\mbox{\small$EI$}\cdot\frac{d}{ds}\mbox{$\kappa$}(s)\!=\!0\vskip-3.61371pt$ (7)
where we canceled the common factor $\mbox{$\kappa$}(s)\!=\!\tfrac{d}{ds}\phi(s)$. Substituting the system equations $\dot{x}(s)\!=\!\cos\phi(s)$ and $\dot{y}(s)\!=\!\sin\phi(s)$ into Eq. (6)–(7) gives
$\mbox{$\lambda$}_{r}\!\cdot\!\left[\begin{array}[]{c c }\cos\phi_{0}&\sin\phi_{0}\\ \sin\phi_{0}&-\cos\phi_{0}\end{array}\right]\left(\begin{array}[]{c }\!\!\dot{x}\mbox{\small$(s)$}\!\!\\ \!\!\dot{y}\mbox{\small$(s)$}\!\!\end{array}\right)=\left(\begin{array}[]{c }\!\!\mbox{$\frac{1}{2}$}\mbox{\small$EI$}\!\cdot\!\mbox{$\kappa$}^{2}(s)+H^{*}\!\!\\ \!\!\mbox{\small$EI$}\!\cdot\!\frac{d}{ds}\mbox{$\kappa$}(s)\!\!\end{array}\right).\vskip-2.8903pt$
Integrating both sides, $x(s)\!=\!\int_{0}^{s}\dot{x}(t)dt$ and $y(s)\!=\!\int_{0}^{s}\dot{y}(t)dt$, gives the cable $(x,y)$ coordinates in terms of its curvature
$\begin{array}[]{cl}&\left(\begin{array}[]{c }\!\!x(s)\!\!\\ \!\!y(s)\!\!\end{array}\right)\!=\!\left(\begin{array}[]{c }\!\!x(0)\!\!\\ \!\!y(0)\!\!\end{array}\right)\\[6.0pt] &+\frac{1}{\mbox{\small$\lambda$}_{r}}\mbox{\small$\left[\begin{array}[]{c c }\cos\phi_{0}&\sin\phi_{0}\\ \sin\phi_{0}&-\cos\phi_{0}\end{array}\right]$}\left(\begin{array}[]{c }\!\!\int_{0}^{s}\mbox{$\frac{1}{2}$}\mbox{\small$EI$}\cdot\mbox{$\kappa$}^{2}(t)dt+H^{*}\!\!\cdot\!s\!\!\\ \!\!\mbox{\small$EI$}\!\cdot\!(\mbox{$\kappa$}(s)\!-\!\mbox{$\kappa$}(\mbox{\small$0$}))\!\!\end{array}\right)\end{array}\vskip-4.33601pt$ (8)
where $s\!\in\![0,L]$. When a flexible cable is steered with equal endpoint tangents, its equilibrium shapes possess inflection points, zero curvature points at which the cable curvature switches sign (Fig. 2). Such cable shapes are called inflectional elastica. [^1] Non-inflectional elastica form spiral-like shapes when held with equal endpoint tangents. Such shapes may be useful in robotic applications, but the current paper focuses on the more common inflectional elastica. The curvature of the inflectional elastica is given by an elliptic cosine function of the cable path length parameter [love][p. 402-404]
$\mbox{$\kappa$}(s)=-\mbox{\small$\lambda$}\mbox{\small$A$}\!\cdot\!\mbox{\sf$cn$}\big(\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}\cdot(s\!+\!s_{0}),\mathrm{k}\big)\hskip 10.00002pt\mbox{\hskip 10.00002pt $s\in[0,L]$}\vskip-5.05942pt$ (9)
where $\mbox{\sf$cn$}(\cdot)$ has an ellipse modulus $0\!<\!\mathrm{k}\!<\!1$ discussed below, $\mbox{$\lambda$}\!=\!\mbox{$\lambda$}_{r}/\mbox{\small$EI$}$, $A$ is an amplitude parameter discussed below, and $s_{0}$ is a phase parameter that is used to locate the cable start point (Fig. 3). Verification that $\mbox{$\kappa$}(s)$ satisfies the energy extremal conditions of Eqs. (3)-(4) appears in the appendix.
Figure 3: The elastica solutions are characterized by the phase parameter, $s_{0}$, the amplitude parameter, $A$, and the modulus parameter, $0\!<\!\mathrm{k}\!<\!1$, determined by $\phi(s^{*})$. Note that $\mbox{\footnotesize$A$}\!=\!2\mathrm{k}/\mbox{\scriptsize$\sqrt{\mbox{$\lambda$}}$}$, where $\lambda$ represents the magnitude of the force applied at the cable endpoints.
The elliptic cosine function is a periodic cosine-like function having two zeroes per period [ellip_book]. The elastica axis shown in Fig. 2 is defined as the line that passes through the elastica inflection points. Some intuition on this axis can be gained by the following observations. In Eq. (8), the term $\mbox{\small$EI$}\!\cdot\!(\mbox{$\kappa$}\mbox{\small$(s)$}\!-\!\mbox{$\kappa$}(\mbox{\small$0$}))$ multiplies the column vector $(\sin\phi_{0},-\!\cos\phi_{0})$. Since $\mbox{$\kappa$}(s)$ is a periodic cosine-like function, $\phi_{0}$ is the elastica axis angle. It also follows from Eq. (8) that the elastica height above its axis is $h(s)\!=\!\mbox{\small$A$}\!\cdot\!\mbox{\sf$cn$}(\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}\!\cdot\!(s\!+\!s_{0}),\mathrm{k})$ (Fig. 3), which explains the amplitude parameter $A$.
The geometric meaning of the modulus parameter, $\mathrm{k}$, is based on the cable tangent, $\phi(s)$, which is given in terms of an elliptic sine function, $\mbox{\sf$sn$}(\cdot)$ (see appendix for verification)
$\sin\big(\tfrac{1}{2}(\phi(s)-\phi_{0})\big)\!=\!-\mathrm{k}\cdot\mbox{\sf$sn$}\big(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot(s+s_{0}),\mathrm{k}\big)\hskip 10.00002pt\mbox{$s\in[0,L]$.}\vskip-5.05942pt$ (10)
The elliptic sine function is a periodic sine-like function satisfying $\mbox{\sf$cn$}^{2}(u)\!+\!\mbox{\sf$sn$}^{2}(u)\!=\!1$. It follows from Eqs. (9)-(10) that $\mbox{\sf$sn$}(\mbox{\scriptsize$\sqrt{\mbox{$\lambda$}}$}\!\cdot\!\mbox{\small$(s\!+\!s_{0})$},\mathrm{k})\!=\!\pm 1$ at the zero curvature points. Substituting this value in Eq. (10) gives
$k=\left|\sin\big(\tfrac{1}{2}(\phi(s^{*})-\phi_{0})\big)\right|\hskip 10.00002pt\mbox{\hskip 10.00002pt $\mbox{$\kappa$}(s^{*})\!=\!0$.}\vskip-5.78172pt$ (11)
The modulus parameter, $0\!<\!\mathrm{k}\!<\!1$, thus represents the flexible cable incidence angle with its elastica axis (Fig. 3). This paper will also use an equivalent parameter that captures the sign of the incidence angle, $1\!<\!\mbox{$\sigma$}\!<\!1$, defined as
$\mbox{$\sigma$}=\cos\big(\phi(s^{*})-\phi_{0}\big)\hskip 10.00002pt\mbox{\hskip 10.00002pt $\mbox{$\kappa$}(s^{*})\!=\!0$,}\vskip-6.50403pt$
where $\mathrm{k}^{2}\!=\!(1\!-\!\mbox{$\sigma$})/2$.
Determination of flexible cable shape from endpoint constraints: Let us formulate three equations that capture the flexible cable equilibrium shapes in terms of the endpoint constraints, $S(0)\!=\!(x(0),y(0),\phi(0))$ and $S(\mbox{\small$L$})\!=\!(x(\mbox{\small$L$}),y(\mbox{\small$L$}),\phi(\mbox{\small$L$}))$. The flexible cable equilibrium shapes will be characterized by three parameters: the co-state $\mbox{$\lambda$}_{r}$ or equivalently $\mbox{$\lambda$}\!=\!\mbox{$\lambda$}_{r}/\mbox{\small$EI$}$, the elastica modulus parameter $\mathrm{k}$ and the phase parameter $s_{0}$.
First consider the parameter $H^{*}$ in Eq. (8). Since $H(s)\!=\!H^{*}$ for $s\!\in\![0,\mbox{\small$L$}]$, the value of $H^{*}$ can be determined at any point along the cable. At the zero curvature points, $\mbox{$\kappa$}(s^{*})\!=\!0$, Eq. (6) gives
$H^{*}\!=\!\mbox{$\lambda$}_{r}\cdot(\cos\phi\mbox{\small$(s^{*})$}\cos\phi_{0}\!+\!\sin\phi\mbox{\small$(s^{*})$}\sin\phi_{0})\!=\!\mbox{$\lambda$}_{r}\mbox{$\sigma$}.\vskip-5.78172pt$
Next consider the amplitude $A$ in Eq. (9). The elastica highest point with respect to its axis occurs at $s\!=\!-s_{0}$, where $s\!+\!s_{0}\!=\!0$ and $\mbox{\sf$cn$}(0,\mathrm{k})\!=\!1$ (Fig. 3). Substituting $H^{*}\!=\!\mbox{$\lambda$}_{r}\mbox{$\sigma$}$ in Eq. (6), then evaluating $H(s)$ at $s\!=\!-s_{0}$ using the relations $\mbox{$\lambda$}\!=\!\mbox{$\lambda$}_{r}/\mbox{\small$EI$}$ and $\mathrm{k}^{2}\!=\!(1\!-\!\mbox{$\sigma$})/2$ gives the amplitude
$\mbox{$\lambda$}_{r}-\mbox{$\frac{1}{2}$}\mbox{\small$EI$}\cdot(\mbox{$\lambda$}\mbox{\small$A$})^{2}\!=\!\mbox{$\lambda$}_{r}\mbox{$\sigma$}\,\,\Rightarrow\,\,\mbox{\small$A$}\!=\!\frac{2\mathrm{k}}{\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}}\,.\vskip-5.78172pt$ (12)
We can now formulate three equations for $\lambda$, $\mathrm{k}$ and $s_{0}$ in terms of the cable endpoint constraints. The cable $(x,y)$ coordinates, Eq. (8), evaluated at $s\!=\!\mbox{\small$L$}$ give two equations
$\begin{array}[]{cl}&\left(\begin{array}[]{c }\!\!x(\mbox{\small$L$})\!\!\\ \!\!y(\mbox{\small$L$})\!\!\end{array}\right)\!-\!\left(\begin{array}[]{c }\!\!x\mbox{\small$(0)$}\!\!\\ \!\!y\mbox{\small$(0)$}\!\!\end{array}\right)\\[6.0pt] &=\frac{1}{\mbox{\small$\lambda$}_{r}}\mbox{\small$\left[\begin{array}[]{c c }\cos\phi_{0}&\sin\phi_{0}\\ \sin\phi_{0}&-\cos\phi_{0}\end{array}\right]$}\left(\begin{array}[]{c }\int_{0}^{L}\mbox{$\frac{1}{2}$}\mbox{\small$EI$}\cdot\mbox{$\kappa$}^{2}(t)dt+(\mbox{\small$\lambda$}_{r}\mbox{$\sigma$})\cdot L\\ \mbox{\small$EI$}\!\cdot\!(\mbox{$\kappa$}(\mbox{\small$L$})\!-\!\mbox{$\kappa$}(\mbox{\small$0$}))\end{array}\right)\\[6.0pt] &=\frac{1}{\mbox{\small$\lambda$}}\mbox{\small$\left[\begin{array}[]{c c }\cos\phi_{0}&\sin\phi_{0}\\ \sin\phi_{0}&-\cos\phi_{0}\end{array}\right]$}\left(\begin{array}[]{c }\int_{0}^{L}\mbox{$\frac{1}{2}$}\mbox{$\kappa$}^{2}(t)dt+(\mbox{\small$\lambda$}\mbox{$\sigma$})\cdot L\\ \mbox{$\kappa$}(\mbox{\small$L$})\!-\!\mbox{$\kappa$}(\mbox{\small$0$})\end{array}\right)\end{array}\vskip-1.4457pt$ (13)
where we substituted $H^{*}\!=\!\mbox{$\lambda$}_{r}\mbox{$\sigma$}$ and $\mbox{$\lambda$}\!=\!\mbox{$\lambda$}_{r}/\mbox{\small$EI$}$. The endpoint curvatures in Eq. (13) are $\mbox{$\kappa$}(0)\!=\!-2\mathrm{k}\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot\mbox{\sf$cn$}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot s_{0},\mathrm{k})$ and $\mbox{$\kappa$}(\mbox{\small$L$})\!=\!-2\mathrm{k}\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\!\cdot\!\mbox{\sf$cn$}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\!\cdot\!(s_{0}\!+\!\mbox{\small$L$}),\mathrm{k})$, based on Eq. (9). The elastica axis direction, $\phi_{0}$, is obtained by evaluating Eq. (10) for $\phi(s)$ at s=0s\!=\!0
$\phi_{0}=\phi(\mbox{\small$0$})+2\sin^{-1}\big(\mathrm{k}\cdot\mbox{\sf$sn$}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\!\cdot\!s_{0},\mathrm{k})\big).\vskip-5.78172pt$ (14)
The third equation is obtained by evaluating Eq. (10) at the cable endpoints s=0s\!=\!0 and $s\!=\!\mbox{\small$L$}$
$\begin{array}[]{cl}&\tfrac{1}{2}\big(\phi(\mbox{\small$L$})-\phi(0)\big)=\\[6.0pt] &\sin^{-1}\big(\mathrm{k}\!\cdot\!\mbox{\sf$sn$}(\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}\!\cdot\!s_{0},\mathrm{k})\big)-\sin^{-1}\big(\mathrm{k}\!\cdot\!\mbox{\sf$sn$}(\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}\!\cdot\!(s_{0}\!+\!L),\mathrm{k})\big).\end{array}\vskip-1.4457pt$ (15)
Equations (13)-(15) provide three equations in $\lambda$, $\mathrm{k}$ and $s_{0}$ whose solution by numerical means gives the flexible cable shape as a function of its relative endpoint states, S(L)S(0)S(L)\!-\!S(0). The cable base frame configuration is defined as $(x(0),y(0),\phi(0))$, and together with S(L)S(0)S(L)\!-\!S(0) gives a total of six configuration variables. This paper describes a steering scheme that maintains equal endpoint tangets, thus giving a total of five configuration variables.
The elastica full period length: The elliptic cosine and sine functions period is $4K(\mathrm{k})$, where $K(\mathrm{k})$ is the complete elliptic integral of the first kind. [^2] It is defined as $K(\mathrm{k})\!=\!\int_{0}^{\pi/2}\!\!\tfrac{d\theta}{\sqrt{1-\mathrm{k}^{2}\sin^{2}\theta}}$, a standard analytic function of the modulus parameter $\mathrm{k}$ [ellip_book]. Denote by $\tilde{L}$ the full period length of the elastica that contains the physical cable. The argument $\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot(s\!+\!s_{0})$ in Eqs. (9)-(10) satisfies the full-period relation $\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot\tilde{L}\!=\!4K(\mathrm{k})$, which gives $\tilde{L}\!=\!4K(\mathrm{k})/\mbox{\small$\sqrt{\mbox{$\lambda$}}$}$.
The 2-D cable steering problem: A flexible cable modeled by Euler’s elastica is held by two robot grippers in a planar environment populated by sparsely spaced obstacles. Find a path for the robot grippers that steers the flexible cable from start to target endpoint states using equal endpoint tangents, such that the cable maintains stable equilibrium shapes while avoiding self-intersection and contact with the environment, except for endpoint contacts at the start or target positions.
The sparse obstacles assumption highlights a caveat in the cable steering problem. Two robot grippers are able to control the flexible cable shape only within the elastica family of solutions. Obstacle avoidance is therefore feasible only among sparsely spaced obstacles (see Section VII).
III Non Self-Intersecting Flexible Cable Shapes
This section uses the elastica solutions to identify the non self-intersecting flexible cable shapes, which remarkably involve only the modulus parameter $\mathrm{k}$.
A flexible cable can possibly self-intersect when its tangent angle with respect to its elastica axis, $\phi(s)\!-\!\phi_{0}$, exceeds $\pm 90^{\circ}$ for some $s\!\in\![0,L]$ (Fig. 3). Using Eq. (10) for $\phi(s)$, the flexible cable can self-intersect only when $\mbox{\small$\left|\sin(\tfrac{1}{2}(\phi\mbox{\small$(s)$}\!-\!\phi_{0}))\right|$}\!\geq\!\tfrac{1}{\sqrt{2}}$ for some $s\!\in\![0,L]$. When this happens, the cable crosses its elastica axis at an angle $\mbox{\small$\left|\sin(\tfrac{1}{2}(\phi\mbox{\small$(s^{*})$}-\phi_{0}))\right|$}\!\geq\!\tfrac{1}{\sqrt{2}}$ where $\mbox{$\kappa$}(s^{*})\!=\!0$. Hence, according to Eq. (11) for the modulus parameter, $0\!<\!\mathrm{k}\!<\!1$, self-intersection can possibly occur when
$\mathrm{k}=\left|\sin\big(\tfrac{1}{2}(\phi(s^{*})-\phi_{0})\big)\right|\geq\mbox{\small$\frac{1}{\sqrt{2}}$}\hskip 10.00002pt\mbox{\hskip 10.00002pt $\mbox{$\kappa$}(s^{*})\!=\!0$.}\vskip-3.61371pt$
Next consider the projection of the flexible cable $(x(s),y(s))$ coordinates on its elastica axis, denoted $\mbox{$\bar{x}$}(s)$. A full period of the elastica that contains the physical cable starts at $s\!+\!s_{0}\!=\!0$ and ends at $s\!+\!s_{0}\!=\!\tilde{L}$, where $\tilde{L}$ is the full-period length of the elastica that contains the physical cable (Fig. 2). Define the equivalent path length parameter $\bar{s}\!=\!s\!+\!s_{0}$. According to Eq. (8)
$\mbox{$\bar{x}$}(\bar{s})=\frac{\mbox{\small$1$}}{\mbox{\small$\lambda$}_{r}}\!\cdot\!\int_{0}^{\bar{s}}\!\!\mbox{$\frac{1}{2}$}\mbox{\small$EI$}\cdot\mbox{$\kappa$}^{2}(t)dt+\frac{\mbox{\small$1$}}{\mbox{\small$\lambda$}_{r}}H^{*}\!\cdot\!\bar{s}\hskip 10.00002pt\mbox{\hskip 5.0pt $\bar{s}\in[0,\tilde{L}]$.}\vskip-5.05942pt$ (16)
Figure 4: Numerical solution of the non self-intersection condition, $\mbox{$\bar{x}$}_{max}\!\leq\!\mbox{$\bar{x}$}(\mbox{\scriptsize$\tilde{L}/2$})$ where $\mbox{$\bar{x}$}_{max}\!=\!\mbox{$\bar{x}$}(\bar{s}_{max})$, gives the modulus parameter $\mathrm{k}_{max}\!=\!0.855$. The range $0\!<\!\mathrm{k}\!<\!\mathrm{k}_{max}$ guarantees that the flexible cable is not self-intersecting.
When $\mathrm{k}\!\geq\!\mbox{\small$\tfrac{1}{\sqrt{2}}$}$, each full period of the elastica contains two fold points located at $\bar{s}_{max}$ and $\tilde{L}\!-\!\bar{s}_{max}$ at which $\mbox{\small$\left|\phi(\bar{s})\!-\!\phi_{0}\right|$}\!=\!90^{\circ}\!$ (Fig. 4). Evaluating this condition using Eq. (10) for $\phi(s)$ gives
$\mathrm{k}\cdot\mbox{\sf$sn$}\big(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot\bar{s}_{max},\mathrm{k}\big)\!=\!\mbox{\small$\frac{1}{\sqrt{2}}$}\,\Rightarrow\,\bar{s}_{max}\!=\!\mbox{\small$\frac{1}{\mbox{\scriptsize$\sqrt{\mbox{$\lambda$}}$}}$}\mbox{\sf$sn$}^{-1}\big(\tfrac{1}{\sqrt{2}\cdot\mathrm{k}},\mathrm{k}\big).\vskip-5.78172pt$
where $\mbox{\sf$sn$}^{-1}(\cdot)$ is the inverse elliptic sine function. As long as $\mbox{$\bar{x}$}(\bar{s}_{max})\leq\mbox{$\bar{x}$}(\tilde{L}/2)$, the fold points do not touch each other (Fig. 4). Hence, using Eq. (16) for $\mbox{$\bar{x}$}(\bar{s})$, the flexible cable is not self-intersecting when
$\int_{0}^{\bar{s}_{max}}\!\!\mbox{$\frac{1}{2}$}\mbox{\small$EI$}\cdot\mbox{$\kappa$}^{2}(t)dt+\mbox{\small$H^{*}$}\!\cdot\!\bar{s}_{max}\leq\int_{0}^{\tilde{L}/2}\!\mbox{$\frac{1}{2}$}\mbox{\small$EI$}\cdot\mbox{$\kappa$}^{2}(t)dt+\mbox{\small$H^{*}$}\!\cdot\!\mbox{\small$\frac{\tilde{L}}{2}$}\vskip-1.4457pt$ (17)
where we canceled the common factor $1/\mbox{\small$\lambda$}_{r}$. Substituting $\mbox{\small$H$}^{*}\!=\!\mbox{$\lambda$}_{r}\!\cdot\!(1\!-\!2\mathrm{k}^{2})$ and $\mbox{$\kappa$}(t)\!=\!-2\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\mathrm{k}\!\cdot\!\mbox{\sf$cn$}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}t,\mathrm{k})$ in Eq. (17) gives the non self-intersection condition
$\begin{array}[]{cl}\int_{0}^{\bar{s}_{max}}\mbox{\small$2$}\mathrm{k}^{2}\hskip 0.80002pt\mbox{\sf$cn$}^{2}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}t,\mathrm{k})dt+(1\!-\!2\mathrm{k}^{2})\!\cdot\!\bar{s}_{max}&\\[4.0pt] \leq\int_{0}^{\tilde{L}/2}\mbox{\small$2$}\mathrm{k}^{2}\hskip 0.80002pt\mbox{\sf$cn$}^{2}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}t,\mathrm{k})dt+(1\!-\!2\mathrm{k}^{2})\!\cdot\!\frac{\tilde{L}}{2}\end{array}\vskip-3.61371pt$
where we canceled the common factor $\mbox{\small$\lambda$}_{r}\!=\!\mbox{\small$EI$}\!\cdot\!\mbox{\small$\lambda$}$. A change of integration variable, $u\!=\!\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}\cdot t$ with $du\!=\!\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}\cdot dt$, gives an equivalent non self-intersection condition that depends only on $\mathrm{k}$
$\begin{array}[]{cl}\int_{0}^{u_{max}(\mathrm{k})}\mbox{\small$2$}\mathrm{k}^{2}\hskip 0.80002pt\mbox{\sf$cn$}^{2}(u,\mathrm{k})du+(1\!-\!2\mathrm{k}^{2})\!\cdot\!u_{max}(\mathrm{k})&\\[4.0pt] \leq\int_{0}^{2K(\mathrm{k})}\mbox{\small$2$}\mathrm{k}^{2}\hskip 0.80002pt\mbox{\sf$cn$}^{2}(u,\mathrm{k})du+(1\!-\!2\mathrm{k}^{2})\!\cdot\!2K(\mathrm{k})\end{array}\vskip-5.78172pt$ (18)
with integration limits $u_{max}(\mathrm{k})\!=\!\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}\!\cdot\!\bar{s}_{max}\!=\!\mbox{\sf$sn$}^{-1}\big(\tfrac{1}{\sqrt{2}\cdot\mathrm{k}},\mathrm{k}\big)$ and $u\!=\!\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}\cdot\tfrac{1}{2}\tilde{L}\!=\!2K(\mathrm{k})$ since $\tilde{L}\!=\!4K(\mathrm{k})/\mbox{\small$\sqrt{\mbox{$\lambda$}}$}$. The two sides of Eq. (18) depend only on $\mathrm{k}$ and intersect at $\mathrm{k}_{max}\!=\!0.855$ (Fig. 4). The range $0\!<\!\mathrm{k}\!<\!\mathrm{k}_{max}$ guarantees that the flexible cable is not self-intersecting, irrespective of the other elastica parameters.
Remark: The range $0\!<\!\mathrm{k}\!<\!0.855$ provides a conservative rule for non self-intersection. Elastica solutions with modulus parameter in the range $0.855\!\leq\!\mathrm{k}\!<\!1$ are self-intersecting, but the physical cable may occupy a short segment of such elastica that is not self-intersecting. $\circ$
IV Flexible Cable Stability
This section uses the elastica solutions to obtain a simple geometric characterization of the stable cable shapes to be used during steering. Recall that the elastica inflection points are zero curvature points (Fig. 2). The following theorem adapted from Sachkov [sachkov07] characterizes the stable cable shapes in terms of the number of inflection points.
Figure 5: (a)-(b) Two families of energy extremal cable shapes, $(x_{1}(s),y_{1}(s))$ determined by $s_{0}\!=\!a_{0}$ and $(x_{2}(s),y_{2}(s))$ determined by $s_{0}\!=\!b_{0}$. All cable shapes have same length of $\mbox{\scriptsize$L$}\!=\!1$. (c) When the two families of extremal cable shapes approach the limit $a_{0}\!=\!b_{0}\!=\!\tfrac{\tilde{L}}{4}$, their common endpoint reaches a conjugate point along the limit cable shape.
Theorem 1 [Stability]: Let a flexible cable be held with fixed endpoint positions and tangents at an equilibrium state. Cable shapes that contain zero, one or two inflection points are possibly stable while cable shapes that contain three or more inflection points are definitely unstable.
[…]
In order to describe the notion of a conjugate point, consider a flexible cable held at a fixed initial point and fixed tangent at this point. The cable equilibrium shape is fully determined as the solution of the cable system equations (Eq. (1)) and the cable adjoint equations (Eq. (3)) for each initial value of the co-states $(\mbox{$\lambda$}_{x},\mbox{$\lambda$}_{y})$ and $\mbox{$\lambda$}_{\phi}(\mbox{\small$0$})$. In Section II, the co-states were replaced by the co-state parameter $\lambda$, the elastica modulus $\mathrm{k}$ and the phase parameter $s_{0}$.
One technique to identify a conjugate point (which marks loss of stability) is described in [gelfand][Definition 27.4]. Consider two families of energy extremal cable shapes sharing a fixed initial point and fixed tangent at this point, parameterized by $\lambda$, $\mathrm{k}$ and $s_{0}$. If the two distinct families merge into a single extremal cable shape at the limit $(\mbox{\small$\lambda$},\mathrm{k},s_{0})\rightarrow(\mbox{\small$\lambda$}^{*},\mathrm{k}^{*},s^{*}_{0})$ while intersecting at their respective endpoints, the endpoint of the limit cable shape associated with $(\mbox{\small$\lambda$}^{*},\mathrm{k}^{*},s^{*}_{0})$ forms a conjugate point along the limit cable shape.
Using this insight, consider two families of energy extremal cable shapes having the same modulus parameter, $\mathrm{k}\!=\!\mathrm{k}^{*}$, the same co-state parameter, $\mbox{\small$\lambda$}\!=\!\mbox{\small$\lambda$}^{*}$, and variable phase parameter $s_{0}$. From Section II, the full period length of the elastica that contains the physical cable is given by $\tilde{L}\!=\!4K(\mathrm{k})/\mbox{\small$\sqrt{\mbox{$\lambda$}}$}$. Hence all cable shapes of both families have full period length of $\tilde{L}\!=\!4K(\mathrm{k}^{*})/\mbox{\small$\sqrt{\mbox{$\lambda$}^{*}}$}$.
Consider the family of energy extremal cable shapes given by Eq. (8), $(x_{1}\mbox{\small$(s)$},y_{1}\mbox{\small$(s)$})$, whose phase parameter is $s_{0}\!=\!a_{0}$
$\left(\begin{array}[]{c }\!\!x_{1}(s)\!\!\\ \!\!y_{1}(s)\!\!\end{array}\right)\!=\!\left(\begin{array}[]{c }\!\!x(0)\!\!\\ \!\!y(0)\!\!\end{array}\right)\!+\!\frac{1}{\mbox{\small$\lambda$}^{*}}\mbox{\small$\left[\begin{array}[]{c c }\cos\phi_{0}&\sin\phi_{0}\\ \sin\phi_{0}&-\cos\phi_{0}\end{array}\right]$}\!\left(\begin{array}[]{c }\!\!\int_{0}^{s}\mbox{$\frac{1}{2}$}\mbox{$\kappa$}^{2}(t)dt+\mbox{$\sigma$}\!\cdot\!s\!\!\\ \!\!\mbox{$\kappa$}_{1}(s)\!-\!\mbox{$\kappa$}_{1}(\mbox{\small$0$})\!\!\end{array}\right)$
where $s\!\in\![0,\mbox{\small$\tilde{L}$}$], $\mbox{$\sigma$}\!=\!1\!-\!2(\mathrm{k}^{*})^{2}$ and $\mbox{$\kappa$}_{1}(s)=-\mbox{\small$\lambda$}^{*}\mbox{\small$A$}\!\cdot\!\mbox{\sf$cn$}(\mbox{\small$\sqrt{\mbox{\small$\lambda$}^{*}}$}\cdot(s\!+\!a_{0}),\mathrm{k}^{*})$ (using Eq. (9)), where $\mbox{\small$A$}\!=\!2\mathrm{k}^{*}/\mbox{\small$\sqrt{\mbox{\small$\lambda$}^{*}}$}$.
Next consider an alternative family of energy extremal cable shapes, $(x_{2}\mbox{\small$(s)$},y_{2}\mbox{\small$(s)$})$, whose phase parameter $s_{0}\!=\!b_{0}$ is set symmetrically with respect to $a_{0}$ about the inflection points: if $a_{0}$ lies in the interval $0,\tfrac{\tilde{L}}{2}]$ then $b_{0}\!=\!\tfrac{\tilde{L}}{2}\!-\!a_{0}$; if $a_{0}$ lies in the complimentary interval $(\tfrac{\tilde{L}}{2},\tilde{L}$] then $b_{0}\!=\!\tfrac{3\tilde{L}}{2}\!-\!a_{0}$. Thus
$\left(\begin{array}[]{c }\!\!x_{2}(s)\!\!\\ \!\!y_{2}(s)\!\!\end{array}\right)\!=\!\left(\begin{array}[]{c }\!\!x(0)\!\!\\ \!\!y(0)\!\!\end{array}\right)\!+\!\frac{1}{\mbox{\small$\lambda$}^{*}}\mbox{\small$\left[\begin{array}[]{c c }\cos\phi_{0}&\sin\phi_{0}\\ \sin\phi_{0}&-\cos\phi_{0}\end{array}\right]$}\!\left(\begin{array}[]{c }\!\!\int_{0}^{s}\mbox{$\frac{1}{2}$}\mbox{$\kappa$}^{2}(t)dt+\mbox{$\sigma$}\!\cdot\!s\!\!\\ \!\!\mbox{$\kappa$}_{2}(s)\!-\!\mbox{$\kappa$}_{2}(\mbox{\small$0$})\!\!\end{array}\right)$
where $s\!\in\![0,\mbox{\small$\tilde{L}$}$], $\sigma$ is the same as above and $\mbox{$\kappa$}_{2}(s)=-\mbox{\small$\lambda$}^{*}\mbox{\small$A$}\!\cdot\!\mbox{\sf$cn$}(\mbox{\small$\sqrt{\mbox{\small$\lambda$}^{*}}$}\cdot(s\!+\!b_{0}),\mathrm{k}^{*})$ with $A$ the same as above.
The two families of cable shapes start at $(x\mbox{\small$(0)$},y\mbox{\small$(0)$})$. One can also verify that the two families start along the same tangent direction, $\phi_{1}(0)\!=\!\phi_{2}(0)$, for each matched choice of $a_{0}$ and $b_{0}$ (Fig. 5(a)-(b)). One can also verify that the two families of cable shapes intersect at their respective endpoints, $)(x_{1}(\mbox{\small$\tilde{L}$}),y_{1}(\mbox{\small$\tilde{L}$}))\!=\!(x_{2}(\mbox{\small$\tilde{L}$}),y_{2}(\mbox{\small$\tilde{L}$}$)), for each matched choice of $a_{0}$ and $b_{0}$ (Fig. 5(a)-(b)).
The key step of the proof is to study the limit of the two families when $a_{0}\!=\!b_{0}$. When $a_{0}\!\in\![0,\tfrac{\tilde{L}}{2}]$, this limit occurs when $a_{0}\!\rightarrow\!\tfrac{\tilde{L}}{4}$ and then $b_{0}\!=\!\tfrac{\tilde{L}}{2}\!-\!a_{0}\!\rightarrow\!\tfrac{\tilde{L}}{4}$ (the elastica 1’st inflection point). When $a_{0}\!\in\!(\tfrac{\tilde{L}}{2},\tilde{L}$], the limit occurs when $a_{0}\!\rightarrow\!\tfrac{3\tilde{L}}{4}$ and then $b_{0}\!=\!\tfrac{3\tilde{L}}{2}\!-\!a_{0}\!\rightarrow\!\tfrac{3\tilde{L}}{4}$ (the elastica 2’nd inflection point). In either case, the limit cable shape starts at an inflection point, passes through a midpoint inflection point and ends after a full period at a third inflection point (Fig. 5(c)). Since the two families of extremal cable shapes intersect at their respective endpoints, the endpoint of the limit cable shape at the third inflection point forms a conjugate point along the limit cable shape. $\square$
Theorem 1 is next applied to flexible cables held with equal endpoint tangents. In this case the stable cable shapes consist of two types ($L$ is the physical cable length, $\tilde{L}$ the full period length of the elastica that contains the physical cable).
Corollary 4,1: Under the conditions of Theorem 1, when a flexible cable is held with equal endpoint tangents, the stable cable shapes either contain a single inflection point at the cable midpoint and then $L\!<\!\tilde{L}$, or contain two interior inflection points and then $L\!=\!\tilde{L}$.
Figure 6: The $(\phi(s),\mbox{$\kappa$}(s))$ contours of $H(s)\!=\!H^{*}$ form convex loops that describe full periods of the elastica. When a flexible cable is held with equal endpoint tangents, the cable occupies a segment with vertically aligned endpoints on a particular loop, with one or two inflection points.
Proof: The possibly stable cable shapes contain zero, one or two inflection points according to Theorem 1. From Section II, the Hamiltonian is constant along energy extremal cable shapes, $H(s)\!=\!H^{*}$ for $s\!\in\![0,L]$. According to Eq. (6) for $H(s)$
$\mbox{$\lambda$}_{r}\cdot\cos(\phi\mbox{\small$(s)$}\!-\!\phi_{0})\!-\!\mbox{$\frac{1}{2}$}\mbox{\small$EI$}\cdot\mbox{$\kappa$}^{2}(s)\!=\!\mbox{$\lambda$}_{r}\cdot(1\!-\!2\mathrm{k}^{2})\hskip 10.00002pt\mbox{$\!\!s\in[0,L]$}\vskip-5.05942pt$ (19)
where we substituted $H^{*}\!=\!\mbox{$\lambda$}_{r}\!\cdot\!(1\!-\!2\mathrm{k}^{2})$. The flexible cable tangent and curvature, $(\phi\mbox{\small$(s)$},\mbox{$\kappa$}\mbox{\small$(s)$})$, are variables while $\mbox{\small$\lambda$}_{r}$, $EI$ and $\mathrm{k}$ are fixed parameters. The $(\phi\mbox{\small$(s)$},\mbox{$\kappa$}\mbox{\small$(s)$})$ contours described by Eq. (19) form convex loops centered at the origin, shown in Fig. 6 for different values of the modulus parameter $\mathrm{k}$ (the other parameters do not affect these loops). Each loop represents full period of the elastica, $s\in[0,\tilde{L}$]. Since each loop crosses the $\phi$-axis exactly at two points at which $\mbox{$\kappa$}(s)\!=\!0$, each full period of the elastica contains exactly two inflection points.
When a flexible cable is held with equal endpoint tangents, its endpoints lie on the same vertical line in the $(\phi(s),\mbox{$\kappa$}(s))$ plane (Fig. 6). The physical cable therefore occupies one of two segments on a particular loop, such that the segment is bounded by vertically aligned endpoints. Any such segment crosses the $\phi$-axis at least once, hence the physical cable must have at least one inflection point. When a segment crosses the $\phi$-axis at a single point (one inflection point), the segment occupies less than a full loop and then $L\!<\!\tilde{L}$ (Fig. 6). When a segment crosses the $\phi$-axis at two points (two inflection points), its vertically aligned endpoints merge and the segment forms a full loop with $L\!=\!\tilde{L}$ (Fig. 6). $\square$
Figure 7: The cable start point is fixed at the origin with tangent fixed along the $x$-axis. The cable distal endpoint varies while its tangent is fixed along the $x$-axis. The outer stability region represents endpoint positions where the stable cable shapes have a midpoint inflection point and $L\!<\!\tilde{L}$. The inner stability region represents endpoint positions where the stable cable shapes have two interior inflection points and $L\!=\!\tilde{L}$.
Determination of stable cable shapes from endpoint positions: The two types of stable cable shapes partition the cable relative endpoint positions as follows. Let the cable start point be fixed at the origin with the cable tangent fixed along the $x$-axis (Fig. 7). The cable relative endpoint space is defined as the position of its distal endpoint, $(x(\mbox{\small$L$}),y(\mbox{\small$L$}))$, with the cable endpoint tangent, $\phi(\mbox{\small$L$})$, fixed along the $x$-axis (Fig. 7).
From Corollary 1, the length of any stable cable shape is either $L\!<\!\tilde{L}$ (one inflection point) or $L\!=\!\tilde{L}$ (two inflection points). As $L$ becomes shorter than $\tilde{L}$, the physical cable occupies smaller portion of the full elastica period. If the modulus parameter is held fixed when $L$ becomes shorter than $\tilde{L}$, the cable flattens while its distal endpoint moves away from the origin. To allow the cable endpoint reach points located on the vertical $y$-axis, a maximal flattening parameter limits the cable flattening to the range $\rho\!\cdot\!\tilde{L}\!\leq\!L\!\leq\!\tilde{L}$, where $0\!<\!\rho\!<\!1$. For instance, $\rho\!=\!1/2$ is used in Fig. 7.
We also need to limit the modulus parameter $\mathrm{k}$ as follows. At $\mathrm{k}\!=\!0$ the cable forms a horizontal straight line with $(x(\mbox{\small$L$}),y(\mbox{\small$L$}))\!=\!(0,\mbox{\small$L$})$ (Fig. 7). As $\mathrm{k}$ increases, the flexible cable folds inward towards the origin. Eventually $\mathrm{k}$ reaches a critical value at which the cable distal endpoint reaches the origin. When $\mbox{\small$L$}\!=\!\mbox{\small$\tilde{L}$}$, this event occurs when the cable forms a figure eight. The critical value, $\mathrm{k}_{c}$, is determined by the condition $\mbox{$\bar{x}$}(\mbox{\small$\tilde{L}$})\!=\!\mbox{$\bar{x}$}(\mbox{\small$0$})$, where $\bar{x}$ is the cable coordinate along its elastica axis. Using Eq. (16) for $\bar{x}$, it can be verified that $\mathrm{k}_{c}\!=\!0.909$. In the following characterization of the stability regions the parameter $\mathrm{k}$ will vary in the interval $0,\mathrm{k}_{c}]$.
Consider the stability regions depicted in relative endpoint space of Fig. 7. The outer boundary consists of two curves, $\alpha$ and $\beta$. The curve $\alpha$ traces the endpoint of a maximally flattened cable satisfying $L\!=\!\rho\cdot\!\tilde{L}$. It is specified by the elastica solution $\mbox{$\alpha$}(\mathrm{k})\!=\!(x(\mbox{\small$L$}),y(\mbox{\small$L$}))$ of Eq. (8), using $\mathrm{k}\!\in\![0,\mathrm{k}_{c}]$ as the curve parameter. This curve consists of an upper piece determined by $s_{0}\!=\!\tfrac{\tilde{L}}{4}\!+\!\tfrac{\tilde{L}-L}{2}$ (cable starts at a distance $\tfrac{\tilde{L}-L}{2}$ from the elastica first inflection point at $\tfrac{\tilde{L}}{4}$), and a lower piece determined by the phase parameter by $s_{0}\!=\!\tfrac{3\tilde{L}}{4}\!+\!\tfrac{\tilde{L}-L}{2}$ (cable starts at a distance $\tfrac{\tilde{L}-L}{2}$ from the elastica second inflection point at $\tfrac{3\tilde{L}}{4}$).
The curve $\beta$ in Fig. 7 traces the cable endpoint when it folds inward towards the origin, obtained by decreasing the elastica full period length from $\mbox{\small$\tilde{L}$}\!=\!\mbox{\small$L$}/\rho$ down to $\mbox{\small$\tilde{L}$}\!=\!\mbox{\small$L$}$ while keeping the modulus parameter fixed at $\mathrm{k}_{c}$. This curve is specified by the elastica solution $)\mbox{$\beta$}(\mbox{\scriptsize$\tilde{L}$})\!=\!(x(\mbox{\scriptsize$\tilde{L}$}),y(\mbox{\scriptsize$\tilde{L}$}$)) of Eq. (8), using $\mbox{\small$\tilde{L}$}\!\in\![\mbox{\small$L$},\mbox{\small$L$}/\rho]$ as the curve parameter. The upper and lower pieces of $\beta$ are determined by the phase parameters $s_{0}\!=\!\tfrac{\tilde{L}}{4}\!+\!\tfrac{\tilde{L}-L}{2}$ and $s_{0}\!=\!\tfrac{3\tilde{L}}{4}\!+\!\tfrac{\tilde{L}-L}{2}$.
The region bounded by $\alpha$ and $\beta$ contains an inner loop marked as $\gamma$ in Fig. 7. The inner loop traces the endpoint of a flexible cable of length $\mbox{\small$L$}\!=\!\mbox{\small$\tilde{L}$}$, having two inflection points at the cable endpoints in addition to a midpoint inflection point. This curve is specified by the elastica solution $\mbox{$\gamma$}(\mathrm{k})\!=\!(x(\mbox{\small$L$}),y(\mbox{\small$L$}))$ of Eq. (8), using $\mathrm{k}\!\in\![0,\mathrm{k}_{c}]$ as the curve parameter. The upper piece of $\gamma$ is determined by the phase parameter $s_{0}\!=\!\tfrac{\tilde{L}}{4}$ (cable starts at the elastica first inflection point and ends after full period length). The lower piece of $\gamma$ is determined by $s_{0}\!=\!\tfrac{3\tilde{L}}{4}$ (cable starts at the elastica second inflection point and ends after full period length).
The relative position of the cable endpoint indicates two types of stable cable shapes when held with equal endpoint tangents. All cable endpoints in the outer region between $\gamma$ and the outer loop formed by $\alpha$ and $\beta$ are associated with stable cable shapes of length $L\!<\!\tilde{L}$. The stable cable shapes for endpoints in this region contain a single inflection point at their midpoint (Fig. 7). All cable endpoints in the inner region bounded by $\gamma$ are associated with stable cable shapes of length $L\!=\!\tilde{L}$. The stable cable shapes for endpoints in this region contain two interior inflection points (Fig. 7).
V Flexible Cable Approximation by Quadratic Arcs
This section describes an approximation of flexible cable stable equilibrium shapes by convex quadratic arcs to be used for efficient collision checking during flexible cable steering.
Figure 8: (a) Approximation of a flexible cable shape (thin curve) having two interior inflection points $p_{1}$ and $p_{3}$ and a maximum curvature point $p_{2}$ by four convex quadratic arcs. (b) Approximation of a flexible cable shape (thin curve) having a midpoint inflection point $p_{2}$ and two extreme curvature points $p_{1}$ and $p_{3}$ by four convex quadratic arcs, two of which form shorter arcs.
Control points selection: Each full period of the elastica consists of four convex arcs having equal length. To describe these arcs, consider the $(\phi\mbox{\small$(s)$},\mbox{$\kappa$}\mbox{\small$(s)$})$ contours depicted in Fig. 6. The contours form concentric convex loops that describe full periods of the elastica, $s\!\in\![0,\mbox{\small$\tilde{L}$}$]. Each loop contains four special points: the upper point $(0,\mbox{$\kappa$}_{max})$ of maximum curvature where $\phi(0)\!=\!0$, the rightmost point $(\phi(\tfrac{\tilde{L}}{4}),0)$ which is the first inflection point at which $\mbox{$\kappa$}\!=\!0$, the bottom point $(0,-\mbox{$\kappa$}_{max})$ of minimum curvature where $\phi(\tfrac{\tilde{L}}{2})\!=\!0$, and the leftmost point $(\phi(\tfrac{3\tilde{L}}{4}),0)$ which is the second inflection point at which $\mbox{$\kappa$}\!=\!0$.
When a flexible cable forms a stable shape with equal endpoint tangents, it occupies up to full period length of the elastica (Corollary 4.1). The cable starts according to its phase parameter, $s_{0}\!\in\![0,\mbox{\small$\tilde{L}$}$], and ends at $s_{0}\!+\!\mbox{\small$L$}$ such that $\mbox{\small$L$}\!\leq\!\mbox{\small$\tilde{L}$}$. The flexible cable control points are its two endpoints, its extreme curvature points when such points exist, and the cable inflection points (one or two such points exist in a full period of the elastica). The flexible cable can thus have up to six control points that divide the cable into up to five convex arcs (Fig. 8).
Bézier polynomial construction: Let the flexible cable control points be indexed according to their appearance on the cable, $p_{0},\ldots,p_{n}$ such that $p_{0}$ and $p_{n}$ are the cable endpoints and $n\!\leq\!5$ (Fig. 8). Every pair of control points, $p_{i}$ and $p_{i+1}$, bounds a convex arc approximated by a convex quadratic arc as follows. The control points $p_{i}$ and $p_{i+1}$ are first augmented with an intermediate control point, $q_{i}$, located at the intersection point of the cable’s tangent lines at $p_{i}$ and $p_{i+1}$ (Fig. 8). The position of $p_{i}$ and $p_{i+1}$ and the tangents at these points are obtained from the elastica solution of Eq. (8). This data feeds a simple formula for computing $q_{i}$ which is omitted. Using Bézier’s technique [bezier], each triplet $p_{i}$, $q_{i}$ and $p_{i+1}$ defines the quadratic polynomial
$P_{i}(t)=(1\!-\!t)^{2}\cdot\vec{p}_{i}+2(1\!-\!t)t\cdot\vec{q}_{i}+t^{2}\cdot\vec{p}_{i+1}\hskip 10.00002pt\mbox{ $t\in[0,1]$}\vskip-7.22743pt$ (20)
where $i\!=\!0\ldots n\!-\!1$. Some intuition for $P_{i}(t)$ can be obtained by considering its midpoint position
$P_{i}(\mbox{\small$\tfrac{1}{2}$})\!=\!\tfrac{1}{4}\vec{p}_{i}+\tfrac{1}{2}\vec{q}_{i}+\tfrac{1}{4}\vec{p}_{i+1}.\vskip-5.78172pt$
The midpoint coefficients form barycentric coordinates that place the midpoint closer to the intermediate control point $q_{i}$ within the triangle formed by $p_{i}$, $q_{i}$ and $p_{i+1}$. Additional intuition can be gained from the quadratic polynomial tangent
$P^{\prime}_{i}(t)=2\big((t\!-\!1)\cdot\vec{p}_{i}+(1\!-\!2t)\cdot\vec{q}_{i}+t\cdot\vec{p}_{i+1}\big)\hskip 10.00002pt\mbox{ $t\in[0,1]$}.\vskip-5.78172pt$
The endpoint tangents $P^{\prime}_{i}(0)\!=\!2(\vec{q}_{i}\!-\!\vec{p}_{i})$ and $P^{\prime}_{i}(1)\!=\!2(\vec{p}_{i+1}\!-\!\vec{q}_{i})$ are aligned with two edges of the triangle formed by $p_{i}$, $q_{i}$ and $p_{i+1}$. In particular, these tangents are collinear with the elastica tangents at $p_{i}$ and $p_{i+1}$ for $i\!=\!0\ldots n\!-\!1$. The Bézier polynomials construction is summarized as Algorithms $1$ and $2$.
Algorithm 1 Control Points Selection
Input: cable start state $(x(0),y(0),\phi(0))$, cable length $L$, cable elastica parameters $\lambda$, $\mathrm{k}$ and $s_{0}$ with $\mbox{\small$\tilde{L}$}\!=\!4K(\mathrm{k})/\mbox{\scriptsize$\sqrt{\mbox{$\lambda$}}$}$.
Data structures: control points set $\cal P$, number of control points 2n+12n\!+\!1, next control point position $s^{*}$.
Initialize: $p_{0}\!=\!(x(0),y(0))$, $\phi_{0}\!=\!\phi(0)$, $\mbox{$\cal P$}\!=\!\{p_{0}\}$, i=0i\!=\!0 , $\mbox{\small$s^{*}$}\!=\!0$.
Set $s^{*}$:
  1. 1.
    If $s_{0}\leq\tfrac{\mbox{\small$\tilde{L}$}}{4}$ and $s_{0}\!+\!\mbox{\small$L$}\geq\tfrac{\mbox{\small$\tilde{L}$}}{4}$ set $\mbox{\small$s^{*}$}=\tfrac{\mbox{\small$\tilde{L}$}}{4}$.
  2. 2.
    If $s_{0}\leq\tfrac{\mbox{\small$\tilde{L}$}}{2}$ and $s_{0}\!+\!\mbox{\small$L$}\geq\tfrac{\mbox{\small$\tilde{L}$}}{2}$ set $\mbox{\small$s^{*}$}=\tfrac{\mbox{\small$\tilde{L}$}}{2}$.
  3. 3.
    If $s_{0}\leq\tfrac{3\mbox{\small$\tilde{L}$}}{2}$ and $s_{0}\!+\!\mbox{\small$L$}\geq\tfrac{3\mbox{\small$\tilde{L}$}}{2}$ set $\mbox{\small$s^{*}$}=\tfrac{3\mbox{\small$\tilde{L}$}}{2}$.
  4. 4.
    If $s_{0}\leq\mbox{\small$\tilde{L}$}$ and $s_{0}\!+\!\mbox{\small$L$}\geq\mbox{\small$\tilde{L}$}$ set $\mbox{\small$s^{*}$}=\mbox{\small$\tilde{L}$}$.
while $s_{0}\leq\mbox{\small$s^{*}$}\leq s_{0}+\mbox{\small$L$}$ do
  $p_{i+1}=(x(\mbox{\small$s^{*}$}\!-\!s_{0}),y(\mbox{\small$s^{*}$}\!-\!s_{0}))$ using elastica solution of Eq. (8).
  $\phi_{i+1}=\phi((\mbox{\small$s^{*}$}\!-\!s_{0})$ using elastica tangent formula of Eq. (10).
  Compute control point $q_{i}$ using $(p_{i},\phi_{i})$ and $(p_{i+1},\phi_{i+1})$.
  $\mbox{$\cal P$}\leftarrow\mbox{$\cal P$}\cup\{q_{i},p_{i+1}\}$.
  $\mbox{\small$s^{*}$}=\mbox{\small$s^{*}$}+\mbox{\small$\tilde{L}$}/4$.
  $i=i+1$.
end while
$n=i$, $p_{n}=(x(\mbox{\small$L$}),y(\mbox{\small$L$}))$, $\phi_{n}\!=\!\phi(0)$.
  Compute control point $q_{n-1}$ using $(p_{n-1},\phi_{n-1})$ and $(p_{n},\phi_{n})$.
$\mbox{$\cal P$}\leftarrow\mbox{$\cal P$}\cup\{q_{n-1},p_{n}\}$.
Output: control points $\mbox{$\cal P$}=\{p_{0},q_{0},p_{1},q_{1},\ldots,p_{n}\}$.
Algorithm 2 Bézier Polynomials Construction
Use Algorithm 11 to compute control points $\mbox{$\cal P$}=\{p_{0},q_{0},p_{1},\ldots,p_{n}\}$, $n\leq 5$.
for $i=0$ to n1n-1 do
Compute quadratic polynomial $P_{i}(t)$ according to Eq. (20).
end for
Output: flexible cable approximation by $P_{0}(t),\ldots,P_{n-1}(t)$.
Example: Fig. 8(a) shows an equilibrium shape that contains two interior inflection points $p_{1}$ and $p_{3}$ and a maximum curvature point $p_{2}$. The elastica parameters are $\mbox{\small$L$}\!=\!\mbox{\small$\tilde{L}$}$, $s_{0}\!=\!0$ and $\mathrm{k}\!=\!0.7746$. The control points divide the cable into four convex arcs of length $\tilde{L}/4$. Fig. 8(b) shows a different stable equilibrium shape that contains a midpoint inflection point $p_{2}$ and two extreme curvature points $p_{1}$ and $p_{3}$. The elastica parameters in Fig. 8(b) are $\mbox{\small$L$}\!=\!\tfrac{2}{3}\mbox{\small$\tilde{L}$}$, $s_{0}\!=\!0.916\mbox{\small$\tilde{L}$}$ and $\mathrm{k}\!=\!0.8515$. The control points divide the cable into four convex arcs with shorter first and fourth arcs (since $\mbox{\small$L$}\!=\!\tfrac{2}{3}\mbox{\small$\tilde{L}$}$). In both examples, the four quadratic polynomials pass through the control points $p_{0},\ldots,p_{4}$ with tangents collinear with the cable tangents at the control points. A measure of the approximation quality is the excess length of the piecewise quadratic approximation relative to the cable length. It is 1.61.6% longer in Fig. 8(a) and 4.24.2% longer in Fig. 8(b). $\circ$
VI Flexible Cable Steering Scheme
This section describes the flexible cable steering scheme which relies on the tools developed in previous sections. The scheme consists of three stages. In the first stage the flexible cable relative endpoint positions are computed as a function of three elastica parameters. The second stage uses the relative endpoint positions and their elastica parameters to initialize a 5-D grid of the cable configuration space. The third stage performs motion planning in the flexible cable 5-D configuration space.
Computation of cable relative endpoint positions: The flexible cable relative endpoint positions, denoted $(\mbox{\small$X_{L}$},\mbox{\small$Y_{L}$})\!=\!(x(\mbox{\small$L$})\!-\!x(\mbox{\small$0$}),y(\mbox{\small$L$})\!-\!y(\mbox{\small$0$}))$, are computed using three elastica parameters: the modulus parameter, $\mathrm{k}$, the phase parameter, $s_{0}$, and the full period length of the elastica, $\mbox{\small$\tilde{L}$}\!=\!4K(\mathrm{k})/\mbox{\small$\sqrt{\mbox{$\lambda$}}$}$. Here $\tilde{L}$ replaces the co-state parameter $\lambda$, since the stable cable shapes described in Section IV are associated with cable lengths $\mbox{\small$L$}\!<\!\mbox{\small$\tilde{L}$}$ and $\mbox{\small$L$}\!=\!\mbox{\small$\tilde{L}$}$.
The three elastica parameters vary in a manner that ensures stable non self-intersecting cable shapes. The modulus parameter varies in the interval $0\!\leq\!\mathrm{k}\!\leq\!\mathrm{k}_{max}$, where $\mathrm{k}\!=\!0$ is a straight-line shape while $\mathrm{k}\!\leq\!\mathrm{k}_{max}$ ensures that the cable is not self-intersecting (Section III). The phase parameter $s_{0}$ varies in the interval $\mbox{\small$L$}/4,3\mbox{\small$L$}/4]$. This interval ensures full coverage of the endpoints $(\mbox{\small$X_{L}$},\mbox{\small$Y_{L}$})$ associated with the stable cable shapes. The parameter $\tilde{L}$ varies in the interval $\mbox{\small$L$},\mbox{\small$L$}/\rho]$, where $0\!<\!\rho\!<\!1$ is a maximal flattening parameter (Section IV).
[…]
$\left(\begin{array}[]{c }\!\!X_{L}(\mathrm{k},s_{0},\mbox{\scriptsize$\tilde{L}$})\!\!\\ \!\!Y_{L}(\mathrm{k},s_{0},\mbox{\scriptsize$\tilde{L}$})\!\!\end{array}\right)\!=\!\frac{\mbox{\small$1$}}{\mbox{\small$\lambda$}}\mbox{\small$\left[\begin{array}[]{c c }\cos\phi_{0}&\sin\phi_{0}\\ \sin\phi_{0}&-\cos\phi_{0}\end{array}\right]$}\left(\begin{array}[]{c }\!\!\int_{0}^{L}\mbox{$\frac{1}{2}$}\mbox{$\kappa$}^{2}(s)ds+\mbox{$\lambda$}\mbox{$\sigma$}\!\!\cdot\!\mbox{\small$L$}\!\!\\ \!\!\mbox{$\kappa$}(\mbox{\small$L$})\!-\!\mbox{$\kappa$}(\mbox{\small$0$})\!\!\end{array}\right)\vskip-1.4457pt$ (21)
where $\phi_{0}$ is specified in Eq. (14), $\mbox{$\sigma$}\!=\!1\!-\!2\mathrm{k}^{2}$, $\mbox{\small$\lambda$}\!=\!\mbox{\small$\lambda$}_{r}/\mbox{\scriptsize$EI$}$, and the cable curvatures $\mbox{$\kappa$}(0)$ and $\mbox{$\kappa$}(\mbox{\small$L$})$ are specified in Eq. (9).
Algorithm 3 Relative Endpoint Positions
[…]
Initialize: $EndPoint\leftarrow[0]$.
/*full period cable shapes (cells 1,2 in Fig. 9)*/
Set $\tilde{L}=L$
for $i=1$ to $n_{\mathrm{k}}$ do
  for $j=1$ to $n_{s_{0}}$ do
[…]
    $s_{0,1}=s_{0}(j)$
    if $0\leq s_{0,1}\!\leq\!\tfrac{L}{2}$ set $s_{0,2}\!=\!\tfrac{L}{2}\!-\!s_{0,1}$ else $s_{0,2}\!=\!\tfrac{3L}{2}\!-\!s_{0,1}$
[…]
  end for
end for
/*less than full period cable shapes (cells 3,4 in Fig. 9)*/
for $i=1$ to $n_{\mathrm{k}}$ do
  for $j=1$ to $n_{\mbox{\scriptsize$\tilde{L}$}}$ do
    Set $s_{0,1}\!=\!\tfrac{3\tilde{L}(j)-2L}{4}$, $s_{0,2}\!=\!\tfrac{5\tilde{L}(j)-2L}{4}$
    Compute $X_{L}\big(\mathrm{k}(i),s_{0,1},\mbox{\scriptsize$\tilde{L}$}(j)\big),Y_{L}\big(\mathrm{k}(i),s_{0,1},\mbox{\scriptsize$\tilde{L}$}(j)\big)$
    $EndPoint[X_{L},Y_{L}]\leftarrow\big(k(i),s_{0,1},\mbox{\scriptsize$\tilde{L}$}(j)\big)$
    Compute $X_{L}\big(\mathrm{k}(i),s_{0,2},\mbox{\scriptsize$\tilde{L}$}(j)\big),Y_{L}\big(\mathrm{k}(i),s_{0,2},\mbox{\scriptsize$\tilde{L}$}(j)\big)$
    $EndPoint[X_{L},Y_{L}]\leftarrow\big(k(i),s_{0,2},\mbox{\scriptsize$\tilde{L}$}(j)\big)$
  end for
end for
Output: EndPointEndPoint array of $(X_{L},Y_{L})$.
Figure 10: (a) The relative endpoint positions $(X_{L},Y_{L})$ computed by Algorithm 3 over the elastica parameter cells shown in Fig. 9. (b) The relative endpoint positions lumped into an $\mbox{\scriptsize$N$}\!\times\mbox{\scriptsize$N$}$ array whose inner black cells represent stable non self-intersecting cable shapes. These cells store the elastica parameters that determine the corresponding cable shape.
[…]
Initialization of cable configuration space grid: The flexible cable 5-D configuration space grid, $CSpace$, consists of $(x(0),y(0),\phi(0))$ and $(\mbox{\small$X_{L}$},\mbox{\small$Y_{L}$})$. The cable base frame position, $(x\mbox{\small$(0)$},y\mbox{\small$(0)$})$, varies in a rectangular environment discretized into $n\!\times\!n$ cells while the cable base frame orientation, $\phi(\mbox{\small$0$})$, is discretized into $m$ cells. The $CSpace$ grid is initialized in two steps. In the first step the relative endpoint positions stored in $EndPoint$ are lumped into an $\mbox{\small$N$}\!\times\!\mbox{\small$N$}$ array, $EndPointGrid$, shown as a $50\!\times\!50$ grid in Fig. 10(b). The inner black cells represent stable non self-intersecting cable shapes. Each of these cells stores two elastica triplets that represent two cable shapes in cells $1$ and $2$ of Fig 9(a), and a single cable shape in cells $3$ and $4$ of Fig 9(a). For each inner black cell of $EndPointGrid$, all cells of $CSpace$ having this $(\mbox{\small$X_{L}$},\mbox{\small$Y_{L}$})$ coordinate are marked as feasible in $CSpace$. These cells correspond to stable and non self-intersecting cable shapes and their elastica parameters are kept in the underlying $EndPointGrid$. All cells of $CSpace$ that project to the outer red cells in Fig. 10(b) are marked as infeasible.
Algorithm 4 Flexible Cable Steering Scheme
Input: cable length $L$; cable CSpaceCSpace grid; start $S$ and target $T$ specified as feasible cells of $CSpace$. Polygonal obstacles $\mbox{$\cal B$}_{1},\ldots,\mbox{$\cal B$}_{k}$.
Data structures: Open list $\cal O$, closed list $\cal C$.
Initialize: $\mbox{$\cal O$}=\{S\}$, $\mbox{$\cal C$}=\emptyset$.
/*Search for shortest $CSpace$ path from start to target*/
while $\mbox{$\cal O$}\neq\emptyset$ do
  Select $z^{*}\!\!\in\!\mbox{$\cal O$}$ according to $A^{*}$ search criterion.
  if $z^{*}=T$ then move to output stage.
  Move $z^{*}$ from $\cal O$ to $\cal C$.
  for each feasible neighbor $z$ of $z^{*}$ in CSpaceCSpace s.t.    $z\mbox{$\not{\!\in}$}\mbox{$\cal C$}$ do
    Extract cable base frame $(x(0),y(0),\phi(0))$ of $z$.
[…]
    if $L=\tilde{L}$ then set $s_{0}=s_{0,1}$ $s_{0,1}=s_{0,2}$ in this case*/
    else set $s_{0}\!=\!s_{0,1}$ when $\kappa(0)>0$ and
        set $s_{0}=s_{0,2}$ when $\kappa(0)<0$.
[…]
    $Q(t)\leftarrow\mathrm{Position}\big(P(t),x(0),y(0),\phi(0)\big)$
    if $Collision(Q(t),\mathcal{B}_{1},\ldots,\mathcal{B}_{k})=\mathrm{FALSE}$
    then add $z$ to $\mathcal{O}$
  end for
end while
Output: CSpaceCSpace path from $S$ to $T$, when such a path exists.
Motion planning in 5-D configuration space: The flexible cable steering scheme is summarized as Algorithm 4. The algorithm is given feasible start and target cells in $CSpace$ and a description of polygonal obstacles in the environment. The main loop of Algorithm 4 selects the current best node in the open list, $z^{*}\!\in\!\mbox{$\cal O$}$, as detailed below. For each feasible neighbor $z\!\in\!CSpace$ of $z^{*}\!$, the elastica parameters of $z$ are extracted and fed into Algorithm 3 that computes the Bézier polynomial $P(t)$ for the cable shape. The cable base frame position extracted from $z$ is used to place the Bézier polynomial in the physical environment as $Q(t)$, then collision of $Q(t)$ with the polygonal obstacles is efficiently checked by $Collision\mbox{\small$(Q(t),\mbox{$\cal B$}_{1},\ldots,\mbox{$\cal B$}_{k})$}$. All collision free neighboring cells of $z^{*}$ are added to the open list, $\cal O$, and the main loop of Algorithm 4 resumes. Eventually the target becomes the best node in $\cal O$ or the open list becomes empty. In the first case the path stored in the closed list, $\cal C$, provides the shortest $CSpace$ path from $S$ to $T$. In the latter case a feasible $CSpace$ path from $S$ to $T$ does not exist.
VII Representative Examples
This section describes execution examples of the steering scheme described in Section VI. The scheme was implemented on MATLAB using elliptic functions library [batista_matlab]. [^3] All simulations run MATLAB-22B on DELL OptiPlex 7080, with Intel I7-10700 CPU running at 2.90 GHz and 16 GB main memory. In all examples the flexible cable maximal flattening parameter is set to $\rho\!=\!0.5$, so that $\tilde{L}$ is at most twice the physical cable length.
Figure 11: Flexible cable steering around a single obstacle. (a) Flexible cable position and shape at the start and target. (b)-(c) Cable folds inward until reaching self-intersection limit while rotating away from the obstacle. (d)-(e) Cable moves to the right along obstacle bottom edge while rotating into horizontal position. (f) Cable stretches outward while circumnavigating the obstacle until its distal endpoint reaches the endpoint target position.
Example 1—single obstacle: The example shown in Fig. 11 steers a unit length flexible cable in the presence of a single obstacle. The cable starts at the base frame position $(x\mbox{\small$(0)$},y\mbox{\small$(0)$},\phi(0))\!=\!(0.12,0.12,-135^{\circ})$ with elastica parameters $(\mathrm{k},s_{0},\mbox{\small$\tilde{L}$})\!=\!(0.671,0,\mbox{\small$L$})$. The start shape forms full period of the elastica with co-axial endpoint tangents (Fig. 11(a)). The cable target base frame position is $(x\mbox{\small$(0)$},y\mbox{\small$(0)$},\phi(0))\!=\!(0,0,180^{\circ})$ with elastica parameters $(\mathrm{k},s_{0},\mbox{\small$\tilde{L}$})\!=\!(0.707,0.9\mbox{\small$L$},1.12\mbox{\small$L$})$. The target shape forms $1/1.12$ of the elastica full period with equal endpoint tangents (Fig. 11(a)).
Algorithm 4 computed the steering path shown in Fig. 11 in $138$ seconds (see video clip). The cable first folds inward until reaching self-intersection limit while rotating away from the obstacle (Fig. 11(b)-(c)). The cable now slides with small clearance along the obstacle bottom edge while rotating into horizontal position (Fig. 11(d)). When the cable distal endpoint reaches the obstacle bottom right corner (Fig. 11(e)), the cable stretches outward by increasing $\tilde{L}$ and decreasing $\mathrm{k}$ until it reaches the endpoint target position (Fig. 11(f)). Note that equal endpoint tangents are maintained during cable steering $\circ$
Figure 12: Flexible cable steering through an opening between two obstacles. (a) Start and target have the same base frame position (see text for distance to target estimates). (b) Cable folds inward while its mid-section slides under left-obstacle bottom edge. (c) Cable reaches the right obstacle. (d)-(e) Cable moves away from both obstacles while stretching. (f) Cable moves between obstacles by further stretching until it reaches the distal endpoint position.
Example 2—two obstacles: The example shown in Fig. 12 steers a unit length flexible cable between two obstacles. The cable start and target share the same base frame position at $(x\mbox{\small$(0)$},y\mbox{\small$(0)$},\phi(0))\!=\!(0.26,0,180^{\circ})$. The cable elastica parameters for the start shape are $(\mathrm{k},s_{0},\mbox{\small$\tilde{L}$})\!=\!(0.5,0,\mbox{\small$L$})$. The start shape forms full period of the elastica with coaxial endpoint tangents (Fig. 12(a)). The cable elastica parameters for the target shape are $(\mathrm{k},s_{0},\mbox{\small$\tilde{L}$})\!=\!(0.707,1.15\mbox{\small$L$},1.32\mbox{\small$L$})$. The target shape forms $1/1.32$ of the elastica full period with equal endpoint tangents (Fig. 12(a)).
Algorithm 4 computed the steering path shown in Fig. 12 in $40$ seconds (see video clip). The cable first folds inward while its mid-section slides with small clearance along the left-obstacle bottom edge (Fig. 12(b)). When the cable mid-section reaches the right obstacle (Fig. 12(c)), it moves away from both obstacles while increasing $\tilde{L}$ in order to stretch and reduce height with respect to its elastica axis (Fig. 12(d)-(e)). The cable distal endpoint now moves between the two obstacles by further increasing $\tilde{L}$ while decreasing $\mathrm{k}$ in order to bring the cable endpoint to its target position (Fig. 12(f)). $\circ$
Practical $A^{*}$ path planning: Algorithm 4 searches the feasible cells of $CSpace$ using the open list nodes, $z\!\in\!\mbox{$\cal O$}$, sorted by trip length estimate from $S$ to $T$ through $z$. Trip length is measured as the sum $l(\mbox{\small$S$},z)\!+\!l(z,\mbox{\small$T$})$, where $l(\mbox{\small$S$},z)$ measures the c-space path length traveled from $S$ to $z$ while $l(z,\mbox{\small$T$})$ estimates the c-space path length yet to be travelled from $z$ to $T$. The simplest choice for $l(z,\mbox{\small$T$})$ would be Euclidean distance to the target computed in CSpaceCSpace
$\begin{array}[]{cl}l(z,\mbox{\small$T$})\!\!&\!=\!\mbox{\small$\left((x_{z}(0)\!-\!x_{T}(0))^{2}\!\!+\!(x_{z}(0)\!-\!x_{T}(0))^{2}\!\!+\!a\!\cdot\!(\phi_{z}(0)\!-\!\phi_{T}(0))^{2}\right.$}\\ &\mbox{\small$\left.+(\mbox{\small$X_{L,z}$}\!-\!\mbox{\small$X_{L,T}$})^{2}\!+\!(\mbox{\small$Y_{L,z}$}\!-\!\mbox{\small$Y_{L,T}$})^{2}\right)^{1/2}$}\end{array}\vskip-5.05942pt$
where $(x_{z}(0),y_{z}(0),\phi_{z}(0),\mbox{\small$X_{L,z}$},\mbox{\small$Y_{L,z}$})$ are the c-space coordinates of $z$, $(x_{T}(0),y_{T}(0),\phi_{T}(0),\mbox{\small$X_{L,T}$},\mbox{\small$Y_{L,T}$})$ are the c-space coordinates of $T$ and a>0a\!>\!0 is a scaling constant. Using the Euclidean distance estimate, Algorithm 4 took $168$ seconds to compute a steering path for Example 2. When the start and target were switched in Example 2, Algorithm 4 took significantly longer time of $816$ seconds to compute a steering path. However, one expects shorter runtime when the cable starts within the opening between the two obstacles.
A practical speedup that also gives shorter computation time when the start and target are switched works when the cable’s start-and-target base frames are located close to each other. In these situations one can compute the shortest collision free path for the cable distal endpoint from start to target (Fig. 12(a)). The shortest collision free path is piecewise linear with vertices located at convex obstacle vertices. Each vertex of the shortest collision free path defines intermediate target position for the cable distal endpoint. The cable is now steered using the Euclidean distance estimate for each intermediate target. In Example 2, three intermediate endpoint targets allowed Algorithm 4 to compute a steering path from start to target in $40$ seconds and in shorter time of $37$ seconds for the switched start and target in this example.
Figure 13: Flexible cable steering in a confined environment containing three obstacles. (a) Flexible cable position and shape at the start and target. (b) Cable folds inward until reaching self-intersection limit while rotating in-place. (c) Cable slides in maximally compressed shape through narrow opening between left obstacles. (d) Cable rotates into vertical position. (e)-(f) Cable stretches outward until both endpoints reach their target positions.
Example 3—multiple obstacles: The example shown in Fig. 13 steers a unit length flexible cable in a confined environment having unit $x$ and $y$ dimensions and containing three obstacles. The cable starts at the base frame position $(x\mbox{\small$(0)$},y\mbox{\small$(0)$},\phi(0))\!=\!(0.03,0.33,0)$ with elastica parameters $(\mathrm{k},s_{0},\mbox{\small$\tilde{L}$})\!=\!(0.675,0.45\mbox{\small$L$},\mbox{\small$L$})$. The start shape forms full period of the elastica with equal endpoint tangents (Fig. 13(a)). The cable target is located at the base frame position $(x\mbox{\small$(0)$},y\mbox{\small$(0)$},\phi(0))\!=\!(0.21,0.12,0)$ with elastica parameters $(\mathrm{k},s_{0},\mbox{\small$\tilde{L}$})\!=\!(0.55,0.5\mbox{\small$L$},1.37\mbox{\small$L$})$. The target shape forms $1/1.37$ of the elastica full period with equal endpoint tangents (Fig. 13(a)).
Algorithm 4 computed the steering path shown in Fig. 13 in $180$ seconds (see video clip). The cable first folds inward until reaching self-intersection limit while rotating in-place into horizontal position (Fig. 13(b)). The cable next slides in maximally compressed shape through the narrow opening between the left obstacles (Fig. 13(c)). When the cable reaches the open area beyond the narrow opening it rotates into vertical position (Fig. 13(d)). The cable now stretches outward at both endpoints by increasing $\tilde{L}$ and decreasing $\mathrm{k}$ (Fig. 13(e)), until both endpoints reach their target positions (Fig. 13(f)). $\circ$
The environment depicted in Fig. 13 highlights the steering scheme limitation to sparsely spaced obstacles. Consider the confined environment of Fig. 14 which has unit $x$ and $y$ dimensions. In Fig. 14(a), the gap $d_{1}\!=\!0.29$ allows cable steering through the opening (Example 3). In Fig. 14(b), the gap $d_{2}\!=\!0.23$ barely allows the cable to move through the opening in a self-folded shape, roughly confined to a rectangle of dimensions $0.15L\!\times\!0.4L$ (Fig. 4). When the gap reduces to $d_{3}\!=\!0.15$, Algorithm 4 reports that a feasible steering path no longer exists between start and target.
Figure 14: Flexible cable steering through an opening between two obstacles. (a) The gap $d_{1}\!=\!0.29$ allows cable steering through the opening. (b) The gap $d_{2}\!=\!0.23$ barely allows the cable to move through the opening in a self-folded shape. (c) The gap $d_{3}\!=\!0.15$ disconnects the start and target when the flexible cable is not allowed to deform against obstacles.
VIII Conclusion
The paper considered flexible cable steering by two robot hands that impose endpoint positions and tangents in two-dimensions. The robot hands steer the flexible cable while maintaining equal endpoint tangents, thus reducing the planning problem into five configuration variables. The paper described the Euler’s elastica solutions for the flexible cable equilibrium shapes. The elastica solutions were then used to develop several tools that were incorporated into a flexible cable steering scheme among sparsely spaced obstacles.
The paper first developed analytic equations that determine the cable elastica parameters in terms of the endpoint positions and tangents imposed by the robot hands. The paper next described the range of the elastica modulus parameter that ensures non self-intersection. Using the notion of conjugate point, the paper established a criterion for the flexible cable stability: all cable shapes that form less than full period of the elastica (midpoint inflection point) or full period of the elastica (two internal inflection points) are automatically stable. Last, the paper described an approximation of the flexible cable equilibrium shapes by convex quadratic arcs that allow efficient collision check against obstacles during steering.
The new tools were incorporated into a steering scheme for flexible cables in two-dimensions by two robot hands. The steering scheme has been implemented and execution examples demonstrated the flexible cable motion between start and target while avoiding self-intersection and collision with sparsely placed obstacles in the environment.
Future research will extend this paper in two important ways. The first extension will consider flexible cable steering by two robot hands under gravity in two-dimensions. The flexible cable equilibrium shapes now minimize the cable elastic and gravitational energies
$\mbox{$\cal E$}=\int_{0}^{L}\!\left(\mbox{$\frac{1}{2}$}\mbox{\small$EI$}\!\cdot\!\mbox{$\kappa$}^{2}(s)+\rho g\!\cdot\!y(s)\right)ds\vskip-5.78172pt$
where $\mbox{$\kappa$}(s)$ is the cable curvature, $y(s)$ is the cable vertical coordinate and $L$ is the cable length. The cable stiffness, $EI$, and the cable mass density, $\rho$, are known parameters such that $m\!=\!\rho\mbox{\small$L$}$ is the cable mass. When $\rho$ is small relative to $EI$, the elastica solutions characterize the cable equilibrium shapes. When $\rho$ is large relative to $EI$, the cable behaves like a chain that attains catenary equilibrium shapes under gravity. A promising approach would be to blend the two analytic solutions into reasonably accurate approximation of the true equilibrium shape. The blending technique will then be augmented with the tools developed in this paper into flexible cable steering scheme under gravity.
A more open ended extension will allow the flexible cable to deform against obstacles during steering from start to target. When friction is sufficiently high at the cable’s contacts with obstacles, the flexible cable will exhibit rolling-like behavior with full segments of the cable contacting obstacles. A promising approach would be to allow the flexible cable to deform against a single wall when steered by two robot hands, while avoiding all other obstacles. Predicting the cable equilibrium shapes will require extension of this paper optimal control formulation to include pure state constraints that capture contacts with obstacles, then use the elastica solutions to model the contact free portions of the cable with endpoints now imposed by the robot hands and the environment. The flexible cable steering scheme will likely require extra state variables that will measure the cable contacts against the environment.
The extensions will eventually allow steering flexible strips by two robot hands in 3-D under the influence of gravity. Suitably designed robot grippers can impose line contacts at the flexible strip opposing ends. By maintaining parallel line contacts during 3-D manipulation, flexible strip steering by two robot hands would become a natural generalization of the steering scheme considered in this paper.
Acknowledgment
Prof. Segev from Ben-Gurion University provided highly appreciated input during the preparation of this paper.
Appendix A: Elliptic Function Details
This appendix verifies the formulas for the flexible cable curvature and tangent given in Section II. Let us first verify that the flexible cable curvature
$\mbox{$\kappa$}(s)\!=\!-2\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}\mathrm{k}\!\cdot\!\mbox{\sf$cn$}(\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}(s_{0}\!+\!s),\mathrm{k})\vskip-5.78172pt$ (22)
solves the elastic energy extremum conditions of Eq. (3)-(4). The two extremum conditions are equivalent to the single condition $H(s)\!=\!H^{*}$ for $s\!\in\![0,\mbox{\small$L$}]$, that can be written as
$\mbox{$\frac{1}{2}$}\mbox{\small$EI$}\cdot\mbox{$\kappa$}^{2}(s)+\mbox{\small$H$}^{*}=\mbox{\small$\lambda$}_{r}\cdot\cos(\phi\mbox{\small$(s)$}\!-\!\phi_{0})\hskip 10.00002pt\mbox{$s\in[0,L]$.}\vskip-5.78172pt$ (23)
Eq. (23) is next convected into an equivalent condition that involves only the cable curvature, $\mbox{$\kappa$}(s)$. Taking the derivative of Eq. (23) with respect to $s$ gives
$\mbox{\small$EI$}\cdot\frac{d}{ds}\mbox{$\kappa$}(s)=-\mbox{\small$\lambda$}_{r}\cdot\sin(\phi\mbox{\small$(s)$}\!-\!\phi_{0})\hskip 10.00002pt\mbox{$s\in[0,L]$.}\vskip-5.78172pt$ (24)
where we used the chain rule and then canceled the common factor $\mbox{$\kappa$}(s)\!=\!\frac{d}{ds}\phi(s)$. Eqs (23)-(24) combine into a single implicit differential equation in $\mbox{$\kappa$}(s)$
$\big(\mbox{$\frac{1}{2}$}\mbox{\small$EI$}\cdot\mbox{$\kappa$}^{2}(s)+\mbox{\small$H$}^{*}\big)^{2}+\big(\mbox{\small$EI$}\cdot\frac{d}{ds}\mbox{$\kappa$}(s)\big)^{2}=\mbox{\small$\lambda$}^{2}_{r}.\vskip-5.78172pt$ (25)
From Section II, $\mbox{\small$H$}^{*}\!=\!\mbox{$\lambda$}_{r}\!\cdot\!(1\!-\!2\mathrm{k}^{2})$. Substituting for $H^{*}$ in Eq. (25) gives
$\big(\mbox{$\frac{1}{2}$}\mbox{\small$EI$}\cdot\mbox{$\kappa$}^{2}(s)+\mbox{\small$\lambda$}_{r}\cdot(1\!-\!2\mathrm{k}^{2})\big)^{2}+\big(\mbox{\small$EI$}\cdot\frac{d}{ds}\mbox{$\kappa$}(s)\big)^{2}=\mbox{\small$\lambda$}^{2}_{r}.\vskip-6.50403pt$ (26)
The derivative of $\mbox{\sf$cn$}(u,\mathrm{k})$ is given by the rule: $\tfrac{d}{du}\mbox{\sf$cn$}(u,\mathrm{k})=-\mbox{\sf$sn$}(u,\mathrm{k})\cdot\mbox{\sf$dn$}(u,k)$, where $\mbox{\sf$sn$}(u,\mathrm{k})$ is the elliptic sine function and $\mbox{\sf$dn$}(u,\mathrm{k})$ is the elliptic delta amplitude [ellip_book]. Applying this derivative rule to $\mbox{$\kappa$}(s)$ in Eq. (22) gives
$\frac{d}{ds}\mbox{$\kappa$}(s)=2\mbox{\small$\lambda$}\cdot\mbox{\sf$sn$}(\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}(s_{0}\!+\!s),\mathrm{k})\cdot\mbox{\sf$dn$}(\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}\!\cdot\!s,\mathrm{k}).\vskip-5.78172pt$ (27)
where we used the chain rule. Substituting for $\mbox{$\kappa$}(s)$ and $\tfrac{d}{ds}\mbox{$\kappa$}(s)$ according to Eqs. (22) and (27) in Eq. (26) gives the equation
$\begin{array}[]{cl}&\big(2\mathrm{k}^{2}\cdot\mbox{\sf$cn$}^{2}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}(s_{0}\!+\!s),\mathrm{k})+(1-2\mathrm{k}^{2})\big)^{2}+\\ &4\mathrm{k}^{2}\cdot\mbox{\sf$sn$}^{2}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}(s_{0}\!+\!s),\mathrm{k})\cdot\mbox{\sf$dn$}^{2}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}(s_{0}\!+\!s),\mathrm{k})=1\end{array}\vskip-5.78172pt$ (28)
where we substituted $\mbox{$\lambda$}\!=\!\mbox{$\lambda$}_{r}/\mbox{\scriptsize$EI$}$ then canceled the resulting common factor $\mbox{$\lambda$}^{2}_{r}$. The first summand in Eq. (28) can be simplified as $(1\!-\!2\mathrm{k}^{2}\mbox{\sf$sn$}^{2}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}(s_{0}\!+\!s),\mathrm{k}))^{2}$. The function $\mbox{\sf$dn$}(u,\mathrm{k})$ satisfies the relation $\mbox{\sf$dn$}^{2}(u,\mathrm{k})\!=\!1\!-\!\mathrm{k}^{2}\cdot\mbox{\sf$sn$}^{2}(u,k)$. Substituting for $\mbox{\sf$dn$}^{2}(u,\mathrm{k})$ in Eq. (28) and using the simplified summand gives the relation
$\begin{array}[]{cl}&1-4\mathrm{k}^{2}\cdot\mbox{\sf$sn$}^{2}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}(s_{0}\!+\!s),\mathrm{k}))\big(1-\mathrm{k}^{2}\mbox{\sf$sn$}^{2}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}(s_{0}\!+\!s),\mathrm{k})\big)\\[4.0pt] &+4\mathrm{k}^{2}\cdot\mbox{\sf$sn$}^{2}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}(s_{0}\!+\!s),\mathrm{k})\big(1-\mathrm{k}^{2}\mbox{\sf$sn$}^{2}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}(s_{0}\!+\!s),\mathrm{k})\big)=1\end{array}\vskip-5.05942pt$
which validates the formula for the flexible cable curvature.
Let us next verify the formula for the flexible cable tangent
$\sin\big(\tfrac{1}{2}(\phi(s)-\phi_{0})\big)=-\mathrm{k}\cdot\mbox{\sf$sn$}\big(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot(s_{0}\!+\!s),\mathrm{k}\big).\vskip-5.05942pt$ (29)
Taking the derivative of both sides of Eq. (29) with respect to $s$ using the derivative rule $\tfrac{d}{du}\mbox{\sf$sn$}(u,\mathrm{k})\!=\!\mbox{\sf$cn$}(u,\mathrm{k})\cdot\mbox{\sf$dn$}(u,k)$ [ellip_book] gives
$\begin{array}[]{cl}&\mbox{$\frac{1}{2}$}\cos(\tfrac{1}{2}\big(\phi(s)-\phi_{0})\big)\cdot\frac{d}{ds}\phi(s)=\\ &-\sqrt{\mbox{$\lambda$}}\mathrm{k}\cdot\mbox{\sf$cn$}\big(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot(s_{0}\!+\!s),\mathrm{k}\big)\cdot\mbox{\sf$dn$}\big(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot(s_{0}\!+\!s),\mathrm{k}\big).\end{array}\vskip-6.50403pt$ (30)
where we used the chain rule. Since $\mbox{$\kappa$}(s)\!=\!\tfrac{d}{ds}\phi(s)$ on the left and $\mbox{$\kappa$}(s)\!=\!-2\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\mathrm{k}\!\cdot\!\mbox{\sf$cn$}(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}(s_{0}\!+\!s),\mathrm{k})$ on the right of Eq. (30), it suffices to validate the relation
$\cos(\tfrac{1}{2}\big(\phi(s)\!-\!\phi_{0})\big)=\mbox{\sf$dn$}\big(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot(s_{0}\!+\!s),\mathrm{k}\big)\vskip-5.05942pt$
or equivalently the relation
$\cos^{2}(\tfrac{1}{2}\big(\phi(s)-\phi_{0})\big)=\mbox{\sf$dn$}^{2}\big(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot(s_{0}\!+\!s),\mathrm{k}\big).\vskip-5.05942pt$
Using the identity $\mbox{\sf$dn$}^{2}(u,\mathrm{k})\!+\!\mathrm{k}^{2}\mbox{\sf$sn$}^{2}(u,\mathrm{k})\!=\!1$ [ellip_book], it suffices to verify the relation
$\cos^{2}(\tfrac{1}{2}\big(\phi(s)\!-\!\phi_{0})\big)=1-\mathrm{k}^{2}\cdot\mbox{\sf$sn$}^{2}\big(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot(s_{0}\!+\!s),\mathrm{k}\big).\vskip-5.78172pt$ (31)
The sum of $\sin^{2}(\tfrac{1}{2}(\phi(s)\!-\!\phi_{0}))$ according to Eq. (29) and $\cos^{2}(\tfrac{1}{2}(\phi(s)\!-\!\phi_{0}))$ according to Eq. (31) becomes
$\begin{array}[]{cl}&\sin^{2}\big(\tfrac{1}{2}(\phi(s)\!-\!\phi_{0})\big)+\cos^{2}\big(\tfrac{1}{2}(\phi(s)\!-\!\phi_{0})\big)=\\[4.0pt] &\mathrm{k}^{2}\!\cdot\!\mbox{\sf$sn$}^{2}\big(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot(s_{0}\!+\!s),\mathrm{k}\big)\!+\!\big(1\!-\!\mathrm{k}^{2}\!\cdot\!\mbox{\sf$sn$}^{2}\big(\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot(s_{0}\!+\!s),\mathrm{k}\big)\big)\!=\!1\end{array}\vskip-5.05942pt$
which validates the formula for the flexible cable tangent.
Appendix B: Self-Intersection of the Inflectional Elastica
In this appendix, we first consider the canonical elastica shape without adding a phase $s_{0}$, translation $(x(0),y(0))$, or rotation $\phi(0)$. That is, an infinite elastica that depends on two parameters, the elliptic modulus parameter $k$ or $\mu$ and the co-state parameter $\lambda$.
We now write the equations of the elastica curve $(x(s),y(s))$ according to the parametrization in [djondjorov]:
$\begin{pmatrix}x(s)\\ y(s)\end{pmatrix}\!=\!\begin{pmatrix}\frac{2}{\sqrt{\lambda}}E\bigg(\mathrm{am}(\sqrt{\lambda}s,\mathrm{k})\bigg)-s\\ A\cdot\mathrm{cn}(\sqrt{\lambda}s,\mathrm{k})\end{pmatrix},\qquad s\in[0,L]$ (32)
where the following relations hold:
$\mathrm{k}=\sqrt{\frac{1-\mu}{2}},\qquad a=\sqrt{\frac{2(1-\mu)}{\lambda}},\qquad\lambda=\left(\frac{4\,\mathrm{K}(\mathrm{k})}{\tilde{L}}\right)^{\!2}.$
In [djondjorov], there is an expression for the height $y_{fold}$, which corresponds to what we referred to as fold points. From this expression, it is possible to derive an expression for $s_{fold}$, the length along the cable where the tangent angle with respect to the elastica axis satisfies $\phi(s_{fold})\pm 90^{\circ}$:
$y_{fold}=\sqrt{-\frac{2\mu}{\lambda}}\ \ \Longrightarrow\ \ s_{\mathrm{fold}}=\frac{1}{\sqrt{\lambda}}\,F\!\Big(\cos^{-1}\!\sqrt{\frac{\mu}{\mu-1}},\,\mathrm{k}\Big).$ (33)
Additionally, we note that the points where the elastica curvature $\kappa(s)$ vanishes are
$s(\kappa=0)=\frac{\tilde{L}}{4}+n\,\frac{\tilde{L}}{2},\qquad n=1,2,\dots.$
We also recall that $\mathrm{k}_{int}$ is the maximum elliptic modulus parameter for which there is no self-intersection (there exist aself-tangency point), $\mathrm{k}_{c}$ is the elliptic modulus parameter corresponding to the figure-eight shape. Thus, we divide the elastica shape into three regions:
Region 1: $\mathrm{k}\in[0,\mathrm{k}_{\max}),$
Region 2: $\mathrm{k}\in[k_{\max},\mathrm{k}_{c}),$
Region 3: $\mathrm{k}\in[\mathrm{k}_{c},1).$
IX Region 1
In this region, there is no self-intersection for any phase $s_{0}$ and any full period $\tilde{L}$, as shown in section III.
X Region 2
For convenience, consider a full period of the elastica in the domain $s\in\left[-\tfrac{\tilde{L}}{2},\tfrac{\tilde{L}}{2}\right]$. We look for the first $s\geq 0$ self-intersection point, denoted as $s_{int}$.
Figure 15: A full period of the elastica in Region 2 of $\mathrm{k}$ values.
From Figure 1, several points can be noted that help us understand the self-intersection point of $k$ values:
  • The first self-intersection point in Region 2 always satisfies $x(s_{int})=0$. Therefore, Eq. (1) equals zero—this is the implicit equation we need to solve to find the intersection point $s_{int}$.
  • In Region 2, the self-intersection point $s_{int}$ will always lie in the interval $\left[\tfrac{\tilde{L}}{4},\tfrac{\tilde{L}}{2}-s_{fold}\right]$.
  • In Region 2, it always holds that $x\!\left(\tfrac{\tilde{L}}{4}\right)>0$ and $x\!\left(\tfrac{\tilde{L}}{2}-s_{fold}\right)<0$, so the function $x(s)$ of Eq. (1) has exactly one root in the interval $\left[\tfrac{\tilde{L}}{4},\tfrac{\tilde{L}}{2}-s_{fold}\right]$. Therefore, the bisection method can be used to find $s_{int}$.
  • In Region 2, the first self-intersection point is always in the negative part of the $y$-axis (Fig. 1). Note that the second self-intersection point always lies in the interval $\left[\tfrac{\tilde{L}}{2}-s_{fold},\tfrac{\tilde{L}}{2}\right]$, but this point is not needed to detect self-intersection.
Additionally, two special cases are worth noting:
  • When $\mathrm{k}=\mathrm{k}_{max}$, it holds that $s_{int}=\tfrac{\tilde{L}}{2}-s_{fold}$.
  • When $\mathrm{k}=\mathrm{k}_{c}$, it holds that $s_{int}=\tfrac{\tilde{L}}{4}$, and the self-intersection point of the canonical elastica occurs at the origin.
After presenting Region 3, we will describe an algorithm that examines all possible cases for self-intersection, based on an analysis I performed, which involved tracking the increase of $s$ and searching for the next self-intersection points.
XI Region 3
In this region as well, we consider a full period of the canonical elastica in the domain $s\in\left[-\tfrac{\tilde{L}}{2},\tfrac{\tilde{L}}{2}\right]$ We look for the first positive self-intersection point, denoted by $s_{int}$.
Figure 16: A full period of the elastica in Region 3 of $\mathrm{k}$ values.
Important note: This paper do not consider Region 3 of $\mathrm{k}$-values for local stability reasons. However, locally stable shapes can exist in Region 3. To determine whether a shape is locally stable, one must check the determinant of the endpoint mapping Jacobian [sachkov_conjugate].
From Figure 2, several points can be observed that help us understand the self-intersection point in Region 3 of $\mathrm{k}$-values:
  • The first self-intersection point in Region 3 always satisfies $x(s_{int})=0$. Therefore, Eq. (1) equals zero—this is the implicit equation we need to solve to find the intersection point $s_{int}$.
  • In Region 3, the self-intersection point will always lie within the interval $s\in\left[s_{fold},\tfrac{\tilde{L}}{4}\right]$.
  • In Region 3, it always holds that $x(s_{fold})>0$ and $x\!\left(\tfrac{\tilde{L}}{4}\right)<0$, so the function $x(s)$ of Eq. (1) has exactly one root in the interval $\left[s_{fold},\tfrac{\tilde{L}}{4}\right]$. Therefore, the bisection method can be used to find $s_{int}$.
  • In Region 3, the self-intersection point always lies in the positive part of the $y$-axis (Fig. 2).
XII Self-Collision Check Algorithm
First, let us specify how we compute $s_{int}$. For this, we use the bisection method, denoted as:
$s_{int}=\text{Bisection}(s_{min},s_{max},\epsilon),$
where $s_{min}$ is the lower bound of the interval where Eq. (1) crosses zero, $s_{max}$ is the upper bound, and $\epsilon$ represents the tolerance for the function to be considered zero.
The following algorithm illustrates the computation of $s_{int}$.
Algorithm 5 GetSint
Input: elastica parameters $\mathrm{k}$, $\tilde{L}$, tolerance $\epsilon$.
Data structures: fold location $s_{\text{fold}}$, intersection $s_{\text{int}}$.
Compute: $s_{\text{fold}}$.
Cases:
  1. 1.
    If $\mathrm{k}\leq\mathrm{k}_{\max}$, set $s_{\text{int}}=\emptyset$.
  2. 2.
    If $\mathrm{k}\in[\mathrm{k}_{max},\,\mathrm{k}_{c}]$, set $s_{\text{int}}=\text{Bisection}\!\big(\tfrac{\tilde{L}}{4},\,\tfrac{\tilde{L}}{2}-s_{\text{fold}},\,\epsilon\big)$.
  3. 3.
    If $\mathrm{k}\in[k_{c},\,1)$, set $s_{\text{int}}=\text{Bisection}\!\big(s_{\text{fold}},\,\tfrac{\tilde{L}}{4},\,\epsilon\big)$.
Output: $s_{\text{int}}$.
[…]
Figure 17: Two full periods of the elastica ($\tilde{L}=L/2$) and the six self-intersection points along the periods for Region 2 of $\mathrm{k}$ values.
For Region 2 of $\mathrm{k}$ values, Figure 3 shows the six possible self-intersection points, where some points satisfy $(x(s_{i}),y(s_{i}))=(x(s_{j}),y(s_{j}))$. Based on this, we perform the self-intersection analysis.
For Region 2 of $\mathrm{k}$ values, the values $s_{1}<s_{2}<s_{3}<s_{4}<s_{5}<s_{6}$ are computed as:
$s_{1}=\tfrac{\tilde{L}}{2}-s_{int},\qquad s_{2}=\tilde{L}-s_{int},\qquad s_{3}=\tfrac{\tilde{L}}{2}+s_{int},$
$s_{4}=\tfrac{3\tilde{L}}{2}-s_{int},\qquad s_{5}=\tilde{L}+s_{int},\qquad s_{6}=\tfrac{3\tilde{L}}{2}+s_{int}.$
From Figure 3, we can see that:
  • If $s_{0}$ starts between 0 and $s_{1}$, the next self-intersection point is $s_{3}$. Therefore, if $s_{0}+L<s_{3}$ there is no self-intersection.
  • If $s_{0}\in(s_{1},s_{2}]$, the next self-intersection point is $s_{5}$. Therefore, if $s_{0}+L<s_{5}$ there is no self-intersection.
[…]
The next figure also shows two elastica periods, but for Region 3 of $\mathrm{k}$ values. In this case, $s_{1},\ldots,s_{6}$ are computed differently.
Figure 18: Two full periods of the elastica ($\tilde{L}=L/2$) and the six self-intersection points along the periods for Region 3 of $\mathrm{k}$ values.
For Region 3 of $\mathrm{k}$ values, the values $s_{1}<s_{2}<s_{3}<s_{4}<s_{5}<s_{6}$ are computed as:
$s_{1}=\tfrac{\tilde{L}}{2}-s_{int},\qquad s_{2}=\tfrac{\tilde{L}}{2}+s_{int},\qquad s_{3}=\tilde{L}-s_{int},\qquad$
$s_{4}=\tilde{L}+s_{int},\qquad s_{5}=\tfrac{3\tilde{L}}{2}-s_{int},\qquad s_{6}=\tfrac{3\tilde{L}}{2}+s_{int}.$
From Figure 4, we can see that:
  • If $s_{0}$ starts between 0 and $s_{1}$, the next self-intersection point is $s_{2}$. Therefore, if $s_{0}+L<s_{2}$ there is no self-intersection.
  • If $s_{0}\in(s_{1},s_{3}]$, the next self-intersection point is $s_{4}$. Therefore, if $s_{0}+L<s_{4}$, there is no self-intersection.
[…]
After expressing these inequality conditions as functions of $s_{int}$ and $\tilde{L}$, we obtain identical expressions and domains for both $\mathrm{k}$ regions. Therefore, Algorithm 6 describes the self-intersection check for all $\mathrm{k}$ values, $0\leq\mathrm{k}<1$, within the inflectional elastica domain (including the figure-eight shape).
Algorithm 6 CheckSelfIntersection
Input: $\mathrm{k}$, $\tilde{L}$, $s_{0}$, cable length $L$, tolerance $\epsilon$.
Compute: $s_{\text{int}}=\text{GetSint}(\mathrm{k},\tilde{L},\epsilon)$.
Output: True if a self–intersection occurs, otherwise False.
Early exit:
  1. 1.
    If $\mathrm{k}\leq\mathrm{k}_{\max}$, return False.
Region 2 & 3: ($\mathrm{k}\in(\mathrm{k}_{\max},\,1)$)
  1. 1.
    If $s_{0}\in\big[0,\,\tfrac{\tilde{L}}{2}-s_{\text{int}}\big]$:
    • If $s_{0}+L<\tfrac{\tilde{L}}{2}+s_{\text{int}}$, return False; else return True.
  2. 2.
    If $s_{0}\in\big(\tfrac{\tilde{L}}{2}-s_{\text{int}},\,\tilde{L}-s_{\text{int}}\big]$:
    • If $s_{0}+L<\tilde{L}+s_{\text{int}}$, return False; else return True.
  3. 3.
[…]
    • If $s_{0}+L<\tfrac{3\tilde{L}}{2}+s_{\text{int}}$, return False; else return True.
XIII Examples
This section presents experiments conducted to evaluate the algorithm for self-intersection detection.
Example 1 demonstrates Region 2 of the $\mathrm{k}$ values. In part (a) of Fig. 5, one can see an elastica shape with parameters $\mathrm{k}=0.88$, $\tilde{L}=L$, $s_{0}=0.23\tilde{L}$. In this case, Algorithm 3 returned FALSE, meaning no self-intersection was detected.
In part (b) of Fig. 5, a small change was made in the elastica parameters so that only $s_{0}$ was modified to $s_{0}=0.2\tilde{L}$. In this case, the algorithm detected a self-intersection.
Figure 19: Example of self-intersection detection by Algorithm 3 in Region 2 of the $\mathrm{k}$ values. Part (a) shows an elastica with parameters $\mathrm{k}=0.88$, $\tilde{L}=L$, $s_{0}=0.23\tilde{L}$, where no self-intersection was detected. Part (b) shows an elastica with parameters $\mathrm{k}=0.88$, $\tilde{L}=L$, $s_{0}=0.2\tilde{L}$, where a self-intersection was detected.
Example 2 demonstrates Region 3 of the $\mathrm{k}$ values. In part (a) of Fig. 6, one can see an elastica shape with parameters $\mathrm{k}=0.99$, $\tilde{L}=2.7L$, $s_{0}=0.75\tilde{L}$. In this case, Algorithm 3 returned FALSE, meaning no self-intersection was detected.
In part (b) of Fig. 6, a small change was made in the elastica parameters so that only $\tilde{L}$ was modified to $\tilde{L}=2.5L$. In this case, the algorithm detected a self-intersection.
Figure 20: Example of self-intersection detection by Algorithm 3 in Region 3 of the $\mathrm{k}$ values. Part (a) shows an elastica with parameters $\mathrm{k}=0.99$, $\tilde{L}=2.7L$, $s_{0}=0.75\tilde{L}$, where no self-intersection was detected. Part (b) shows an elastica with parameters $\mathrm{k}=0.99$, $\tilde{L}=2.5L$, $s_{0}=0.75\tilde{L}$, where a self-intersection was detected.
Appendix C: Static Elastica Experiments
This appendix presents static-shape experiments that compare the planar Euler elastica solutions used in the paper with the shape of a real flexible cable. The purpose is to experimentally validate that the elastica model provides an accurate description of the cable shape during quasi-static manipulation.
Figure 21: Static shapes of the cable tie and corresponding elastica solutions. Parts (a)–(f) show the measured cable-tie shapes in the horizontal plane. Parts (g)–(l) show the corresponding planar elastica solutions. In the elastica plots, the black curve represents the elastica, blue markers denote sampling points on the elastica, red markers denote the measured cable-tie points, and the red circle marks the cylindrical obstacle used in some of the shapes.
XIII-A Experimental Setup
The experiments were conducted with a plastic cable tie (zip tie) of length $L_{\mathrm{zip}}=470\,\mathrm{mm}$, width $w_{\mathrm{zip}}=9\,\mathrm{mm}$, and thickness $t_{\mathrm{zip}}=1\,\mathrm{mm}$. One end was rigidly clamped, while the other end was grasped by a Robotiq 2F-85 parallel gripper mounted on a UR5e manipulator.
The cable tie was manipulated in a horizontal plane for static-shape acquisition. A millimeter-grid paper sheet was placed beneath the tie for measurement, and the coordinates of marked points along the tie were manually recorded.
Fifteen approximately uniform arc-length measurement points were defined along the tie. In practice, points $i=2,\dots,14$ were measured due to limited access near the clamped and grasped endpoints.
XIII-B Measurement Errors
Three primary error sources affect the measurements:
  • Grid and thickness error: Due to the finite width of the millimeter-grid paper and the cable-tie thickness, each measured point has an inherent uncertainty of approximately $t_{\mathrm{zip}}/2$ in the direction normal to the elastica.
  • Arc-length projection error: Manual projection of points onto the grid introduces small errors $\Delta s_{i}$ along the curve tangent, with an unknown sign.
  • Systematic errors: Residual curvature of the tie, surface unevenness, and accumulated plastic deformation may introduce additional biases, but all remained small relative to the overall shape.
A preliminary error measurement was performed on a fully stretched zip-tie. The average markers measurement error was determined to be a disc of radius 1.2 mm.
XIII-C Tested Shapes and Notation
Six planar shapes were recorded, denoted $\mathrm{shape}\ 1,\dots,6$. For each shape $j$, the elastica and measured coordinates of point $i$ are denoted respectively as:
$\mathbf{p}^{\mathrm{e}}_{i,j}=(x^{\mathrm{e}}_{i,j},y^{\mathrm{e}}_{i,j}),\qquad\mathbf{p}^{\mathrm{z}}_{i,j}=(x^{\mathrm{z}}_{i,j},y^{\mathrm{z}}_{i,j}).$
The clamped endpoint is $\mathbf{p}^{\mathrm{e}}_{0,j}=(0,0)$ for all shapes.
Table I: Elastica parameters and end points for the six static shapes.
Shape $\mathrm{k}$ $\tilde{L}$ $s_{0}$ $x(L)[mm]$ $y(L)[mm]$
1 0.4637 $L_{zip}$ 0.2350 365.87 0
2 0.6021 $L_{zip}$ 0.2193 278.81 79.99
3 0.6743 $L_{zip}$ 0.2135 217.25 102.37
4 0.7640 $L_{zip}$ 0.2135 143.21 83.76
5 0.7917 $L_{zip}$ 0.2272 136.59 28.75
6 0.7395 $L_{zip}$ 0.2350 187.75 0
Table II: Elastica coordinates $\mathbf{p}^{\mathrm{e}}_{i,j}$ ($m$).
$i$ Shape 1 Shape 2 Shape 3 Shape 4 Shape 5 Shape 6
1 (28.79, 5.58) (28.16, 7.57) (27.74, 8.61) (26.98,10.20) (26.60,10.95) (27.22, 9.80)
2 (54.73,21.36) (49.60,28.82) (46.16,32.36) (40.68,36.74) (39.51,37.96) (44.15,34.70)
3 (75.64,43.49) (60.22,57.23) (50.24,62.33) (36.38,66.59) (36.33,68.06) (48.62,64.72)
4 (93.18,67.83) (62.87,87.08) (44.07,91.62) (21.08,92.29) (24.68,95.66) (46.66,94.65)
5 (110.75,91.51) (63.09,116.58) (34.25,119.43) ( 2.40,115.11) (12.11,122.34) (45.38,124.09)
6 (132.50,112.80) (66.63,146.81) (26.46,148.87) (-14.66,140.32) ( 5.13,151.88) (51.75,153.73)
7 (158.37,126.69) (78.23,173.75) (27.38,178.18) (-22.29,168.57) (11.67,180.24) (70.29,176.17)
8 (188.43,130.53) (100.88,193.69) (42.03,204.47) (-13.55,197.20) (35.09,198.57) (99.36,182.96)
9 (217.67,122.52) (130.47,199.37) (68.96,217.67) (11.87,212.70) (64.55,194.67) (125.87,169.11)
10 (242.15,105.38) (158.67,189.98) (98.24,213.27) (40.85,207.50) (84.20,172.63) (139.75,142.88)
11 (262.23, 82.47) (181.47,169.91) (122.56,195.21) (62.91,186.85) (91.14,143.12) (142.39,112.60)
12 (279.26, 58.39) (198.52,145.87) (139.94,171.43) (76.34,160.65) (90.80,113.64) (140.05, 83.19)
13 (297.92, 34.29) (215.06,120.25) (155.64,145.29) (87.42,132.24) (90.70, 83.16) (139.72, 52.74)
14 (320.73, 14.16) (235.19, 97.45) (174.48,121.39) (102.40,105.78) (98.53, 53.89) (148.60, 23.81)
15 (348.04, 2.11) (261.09, 82.72) (199.60,105.43) (125.70, 87.38) (119.24, 32.88) (170.29, 3.78)
Table III: Measured cable-tie coordinates $\mathbf{p}^{\mathrm{z}}_{i,j}$ ($m$).
$i$ Shape 1 Shape 2 Shape 3 Shape 4 Shape 5 Shape 6
2 (53.82,19.62) (46.93,28.24) (44.97,32.32) (39.50,36.93) (37.50,38.44) (42.47,34.34)
3 (73.89,42.09) (57.99,55.90) (47.50,62.04) (32.66,66.69) (32.98,68.15) (46.50,64.50)
4 (91.91,65.21) (61.50,85.49) (41.48,91.65) (17.90,93.30) (22.45,96.22) (46.00,95.05)
5 (108.88,89.68) (62.00,115.99) (33.97,119.66) ( 0.39,116.81) (11.96,123.19) (45.50,124.48)
6 (130.80,111.10) (64.99,145.88) (26.49,149.07) (-15.05,142.72) ( 6.50,153.02) (51.46,153.80)
7 (156.66,126.53) (76.91,173.22) (27.99,178.38) (-22.50,171.01) (12.93,181.25) (68.77,176.58)
8 (186.46,130.50) (99.22,193.05) (42.35,205.14) (-14.10,199.20) (35.13,200.27) (97.93,184.50)
9 (215.78,123.55) (127.96,199.50) (67.57,218.00) (12.08,215.51) (64.25,197.57) (124.64,171.66)
10 (239.66,106.13) (156.74,190.07) (97.30,213.04) (40.76,210.06) (84.55,174.79) (139.52,146.38)
11 (260.60,83.70) (180.12,171.18) (121.33,195.66) (62.59,188.72) (92.00,145.97) (143.50,116.52)
12 (278.59,60.21) (197.08,147.23) (140.07,172.74) (77.04,163.31) (92.50,116.52) (142.00, 86.04)
13 (297.13,36.17) (213.59,121.21) (156.58,146.73) (88.54,134.80) (92.50, 85.96) (142.00, 54.95)
14 (319.72,15.58) (234.67, 99.12) (175.15,122.14) (102.61,108.19) (98.56, 55.77) (148.07, 20.25)
XIII-D Error Metrics
The absolute error of point $i$ in shape $j$ is
$e_{i,j}=\bigl\|\mathbf{p}^{\mathrm{e}}_{i,j}-\mathbf{p}^{\mathrm{z}}_{i,j}\bigr\|.$
Following the measurement protocol in the experimental document, the relative error is defined as
$\varepsilon_{i,j}=100\,\frac{\big|\|\mathbf{p}^{\mathrm{e}}_{i,j}\|-\|\mathbf{p}^{\mathrm{z}}_{i,j}\|\big|}{\|\mathbf{p}^{\mathrm{e}}_{i,j}-\mathbf{p}^{\mathrm{e}}_{0,j}\|},$
that is, the absolute deviation divided by the Euclidean distance from the clamped endpoint.
Table IV: Absolute errors $e_{i,j}$ ($m$).
$i$ Shape 1 Shape 2 Shape 3 Shape 4 Shape 5 Shape 6
2 1.96 2.73 1.19 1.20 2.06 1.71
3 2.24 2.59 2.75 3.72 3.35 2.13
4 2.90 2.09 2.59 3.34 2.29 0.77
5 2.61 1.24 0.36 2.63 0.85 0.40
6 2.39 1.88 0.20 2.43 1.77 0.30
7 1.72 1.42 0.64 2.45 1.61 1.57
8 1.96 1.77 0.74 2.07 1.69 2.10
9 2.14 2.50 1.43 2.81 2.90 2.83
10 2.60 1.93 0.96 2.56 2.18 3.50
11 2.03 1.84 1.30 1.89 2.97 4.07
12 1.94 1.98 1.31 2.75 3.35 3.45
13 2.04 1.75 1.72 2.79 3.32 3.17
14 1.74 1.75 1.00 2.42 1.87 3.59
Mean 2.17 1.96 1.24 2.54 2.39 2.27
Std. dev. 0.35 0.43 0.75 0.62 0.79 1.27
Table V: Relative errors $\varepsilon_{i,j}$ (%).
$i$ Shape 1 Shape 2 Shape 3 Shape 4 Shape 5 Shape 6
2 2.49 4.52 1.77 1.35 1.99 2.73
3 2.53 3.04 2.40 2.14 1.86 1.77
4 2.21 1.94 1.04 0.36 0.01 0.06
5 1.81 0.78 0.11 1.45 0.66 0.30
6 1.37 0.94 0.13 1.72 0.78 0.02
7 0.70 0.54 0.16 1.44 0.60 0.08
8 0.70 0.60 0.34 1.03 0.83 0.32
9 0.45 0.52 0.04 1.31 1.29 0.62
10 0.74 0.46 0.25 1.17 1.08 1.17
11 0.42 0.04 0.11 0.84 1.68 1.83
12 0.09 0.14 0.49 1.52 2.28 1.92
13 0.19 0.32 0.78 1.73 2.61 1.95
14 0.29 0.06 0.46 1.28 0.83 0.69
Mean 1.08 1.07 0.63 1.34 1.27 1.04
Std. dev. 0.89 1.33 0.72 0.44 0.76 0.91
XIII-E Discussion
Across all six shapes, the absolute and relative errors remain small. Absolute deviations are typically a few millimeters, while relative errors mostly remain below $2\%$. These results validate the suitability of Euler’s elastica as an accurate model for the planar equilibrium shapes of the flexible cable used in this work.